### v3

# Towards Automated Identification of Violation Symptoms of Architecture Erosion

Abstract
Architecture erosion has a detrimental effect on maintenance and evolution, as the implementation drifts away from the intended architecture. To prevent this, development teams need to understand early enough the symptoms of erosion, and particularly violations of the intended architecture. One way to achieve this, is through the automated identification of architecture violations from textual artifacts, and particularly code reviews. In this paper, we developed 15 machine learning-based and 4 deep learning-based classifiers with three pre-trained word embeddings to identify violation symptoms of architecture erosion from developer discussions in code reviews. Specifically, we looked at code review comments from four large open-source projects from the OpenStack (Nova and Neutron) and Qt (Qt Base and Qt Creator) communities. We then conducted a survey and semi-structured interviews to acquire feedback from the involved participants who discussed architecture violations in code reviews, to validate the usefulness of our trained classifiers. The results show that the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100 and 300-dimensional models. In addition, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperforms the individual classifiers. Finally, the findings derived from the online survey and interviews conducted with the involved developers reveal that the violation symptoms identified by our approaches have practical value and can provide early warnings for impending architecture erosion.
Index Terms:
Architecture Violation, Architecture Erosion Symptom, Natural Language Processing, Machine Learning, Deep Learning
1 Introduction
Software architecture is regarded as critical in developing large and complex systems, as it reflects the system structure and behavior and acts as a bridge between business goals and the implemented system [1]. However, the phenomenon of architecture erosion has become one of the major challenges of architecting, impeding software maintenance and undermining software sustainability [2, 3, 4]. Architecture erosion refers to the increasing and accumulating divergences between implemented and intended architecture [2, 5]. It is often described by various terms in the literature [2] and practice [6], such as architectural decay, degradation, and deterioration.
As evidenced in previous studies [6, 2, 7, 8, 9], various symptoms or signs can be observed when architecture erosion occurs during a software system’s life cycle. Our recent systematic mapping study [2] reported four categories of such architecture erosion symptoms: structural symptoms (e.g., cyclic dependencies), violation symptoms (e.g., layering violation), quality symptoms (e.g., high defect rate), and evolution symptoms (e.g., extensive ripple effects of changes).
In this study, we specifically concentrate on violation symptoms, as those are the most immediate symptoms of architecture erosion and they require the most attention from practitioners [10, 11]. This symptom type stems from the definition of architecture erosion [2], as a violation causes the implemented architecture to diverge from the intended one. Architecturally-relevant violations in software systems include violations of design principles, architecture patterns, decisions, and requirements [2]. The accumulation of violation symptoms can render the architecture completely untenable [12]. For brevity, we refer to violation symptoms of architecture erosion as violation symptoms throughout the remainder of this paper.
While a few temporary violation symptoms might be innocuous and might not result in software errors, the accumulation of architecture violations can affect architecture erosion [5, 12, 13]; consequently, erosion negatively affects quality attributes, including maintainability and performance [6, 13]. Thus, it is essential to identify and monitor violation symptoms to facilitate the maintenance of the system architecture and further reveal inconsistencies between the implemented and the intended architecture. This is the first step towards, eventually, repairing or mitigating architecture erosion [12, 2].
Violation symptoms can be identified through source code analysis, but that has certain limitations. As pointed out in previous studies [14, 2, 11], a number of factors hinder the effective utilization of static code analysis tools in precisely identifying violation symptoms through source code. For example, such tools only detect a limited set of types of architecture violations, they only support few programming languages, and they do not capture the semantics of violation symptoms as they rely solely on code metrics. Alternatively, violations symptoms can be manually identified by analyzing textual artifacts that contain information related to system architecture and design, such as code comments, code reviews, and issue trackers [10]. In this context, recent work [10, 11] has introduced a taxonomy on violation symptoms, that can be used as a checklist to detect such symptoms during code review. However, it could be tedious, time-consuming, and potentially error-prone to manually identify potential architecture violations [11]. Automated techniques could address this problem, but to the best of our knowledge, no prior study has investigated the automated identification of violation symptoms in textual artifacts, such as code review comments.
To address this gap, we conduct an exploratory study to assess the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. The automated identification of violation symptoms could act in a complementary way to modern code review. Modern code review usually addresses code changes in general through knowledge sharing [15], while automated identification of violation symptoms can offer an extra layer of scrutiny. For example, such an automated approach can issue warnings for potential or ignored violation symptoms, preventing the integration of code changes with architectural violations into the code base, and ensuring a more robust and reliable development process. Moreover, automatically identifying violation symptoms from code review comments could offer valuable insights at the architecture level and facilitate the corresponding knowledge sharing between developers and reviewers. The ultimate purpose of our approach is to complement the modern code review process, through warning of and guarding against architectural violation issues, eventually maintaining architectural integrity.
In this work, we developed 15 classifiers based on Machine Learning (ML) and 4 classifiers based on Deep Learning (DL), to identify violation symptoms from developer discussions in code reviews of four large open-source software projects from the OpenStack and Qt communities. Our results show that the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803. We also found out that utilizing fastText pre-trained word embedding can achieve relatively good results and it helps to improve the classifiers’ performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100 and 300-dimensional models. Finally, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperform the individual classifiers.
We then conducted an online survey and semi-structured interviews to validate the effectiveness of automated identification of violation symptoms in practice. We received responses from 24 survey participants and 6 interviewees, all of whom were involved in the discussion of the collected review comments concerning architecture violations. The results of our survey and interview show that the practitioners’ perception of the classifiers’ results confirms the promising applications of automated identification of violation symptoms in practice. Especially, such classifiers can enable and inspire the participants to find, prioritize, and handle architecturally-related violation issues.
The main differences between our work and previous studies that focused on utilizing tools to identify architectural issues by source code analysis are that: (1) our work focuses on exploring the identification of violation symptoms through textual artifacts, which reflects more closely the practitioners’ perspective on such symptoms, compared to source code analysis; (2) our work explored the possibility of identifying violation symptoms of architecture erosion by using ensemble learning techniques, which bring advantages over single ML/DL algorithms. The main contributions of this work are the following:
  • This is the first work to explore the application of ML and DL-based classifiers regarding the automated identification of violation symptoms in code review comments.
  • We compared the performance of several trained classifiers using popular ML and DL-based algorithms, as well as the impact of three pre-trained word embeddings on classification performance.
  • We gained insights from participants regarding the usefulness of ML and DL-based classifiers in identifying architecture violations during code review.
The remainder of the paper is organized as follows. Section 2 introduces background information. Section 3 describes the research questions, study design, and experimental setup. The answers to the research questions are presented in Section 4. We discuss the implications of the study results in Section 5 and the threats to validity in Section 6. Section 7 reviews related work, and Section 8 concludes this work with future directions.
2 Background
In this section, we introduce a) the background of architecture erosion and its symptoms; b) architecture conformance checking, which is the most commonly used approach to detect architecture violations; and c) the modern code review process in Gerrit [^1] https://www.gerritcodereview.com/, as this is used in our dataset.
2.1 Architecture Erosion and Related Symptoms
In the past decades, a wide variety of studies are concerned with the architecture erosion phenomenon, which has been extensively discussed and described with various terms [6, 2], such as architecture decay [16, 17], degradation [18, 19], degeneration [20].
Architecture erosion manifests in various symptoms during development and maintenance. A symptom is a partial sign or indicator of the emergence of architecture erosion [3, 2, 10]. According to a recent mapping study [2], erosion symptoms can be classified into four categories (i.e., structural symptoms, violation symptoms, quality symptoms, and evolution symptoms).
Previous studies investigated different symptoms of architecture erosion. Mair et al. [21] proposed a formalization method regarding the process of repairing eroded architecture by finding violation symptoms and recommending optimal repair sequences. Le et al. [3, 17] regarded architectural smells as structural symptoms and they provided metrics to detect instances of architecture erosion by analyzing the detected smells. Martin [22] deemed that the evolution symptoms of architecture erosion include rigidity (a tendency to resist changes), fragility (a tendency to break frequently when modifications occur), immobility (inability to be reused), and viscosity (reduced effectiveness and efficiency due to design or environment issues).
TABLE I: Examples of violation symptoms of architecture erosion collected from Nova, Neutron, Qt Base, and Qt Creator
Project Example of violation symptoms of architecture erosion
Nova But here don’t we have to make a upcall from compute to api db, which will violate api/cell isolation rules. Is there any workaround in this case?
We could get some race conditions when starting the scheduler where it would not know the allocation ratios and would have to call the computes, which is a layer isolation violation to me.”
Neutron Having said all of that: I get that I’m violating an abstraction layer in LinkLocalAddressPair and that this is surprising (and therefore bad).
As you pointed out, since this patch is about the router availability zone it seems like a layer violation doing it in the mech driver.
Qt Base This one seems to be a layering violation: why does QNetworkRequest understand QHttpNetworkConnection?
This breaks the abstraction layer of publicQtConfig and is annoyingly verbose. instead, publicQtConfig should have an optional parameter ’name’ to override the built-in ‘feature’ member.
Qt Creator The unit tests are designed to not depend on any qt creator library to make the dependency breaking easier. Sometimes it works but it is not designed that way.
This looks like we break the design. Why do we can not hold a pointer to the plugin?
As mentioned in the Introduction section, we focus on the most direct erosion symptoms of architecture erosion (i.e., violation symptoms). Table I presents several examples of violation symptoms of architecture erosion from the used dataset (see Section 3); a detailed comparison of our work and related work is given in Section 7.
2.2 Architecture Conformance Checking
Architecture Conformance Checking (ACC) techniques are the most commonly-used approaches to detect architecture violations [2]. They either perform static or dynamic analysis, by comparing the structure of the intended architecture (provided by the architects) with the extracted architecture information from source code. For example, Miranda et al. [23] presented an ACC approach based on static code analysis techniques with a supporting tool ArchRuby. Their tool provides means to control the erosion process by reporting and visualizing architectural violations in two high-level architectural models, namely reflexion models and Dependency Structure Matrices (DSMs).
In addition, rule-based conformance checking approaches are also employed to identify architecture violations. For example, previous studies identified architecture violations by checking explicitly defined architectural rules [13, 24, 25]. Moreover, it is viable to check architecture conformance and identify architecture violations by defining and describing the systems through Architecture Description Languages (ADLs) [25, 26], or Domain-Specific Languages (DSLs) [27, 28].
Despite their potential, the previously mentioned approaches for identifying architecture violations have clear limitations. For instance, much effort is required to address the challenges of understanding the architecture design (e.g., concepts and relations), defining architectural rules (or description languages) in advance, and establishing a mapping between architectural elements and source code [2, 11]. Additionally, other limitations, such as lack of generalizability, lack of visualization of architectural views, and insufficient tooling support, have impeded the wide adoption of these approaches in practice [2].
During software development, textual artifacts (e.g., code review comments) contain rich and valuable information regarding architectural design and changes apart from source code. Mining architectural violations from textual artifacts can significantly facilitate code comprehension for software engineers [11]. In contrast to the previously mentioned approaches, our study focuses specifically on mining violation symptoms from textual artifacts.
2.3 Code Review in Gerrit
Code review is a crucial software development activity that involves the systematic examination of assigned code to identify defects and improve software quality [29]. A methodical code review process not only enhances the quality of software systems, but also facilitates the sharing of development knowledge and prevents the release of unstable and defective products. Over time, code review practices have become increasingly important, have been widely adopted in modern software development, and have received extensive tool support. In fact, tool-based code review has become the norm in both industry and open-source communities, with a variety of available code review tools, such as Me’s Phabricator [^2] https://www.phacility.com/, VMware’s Review-Board [^3] https://www.reviewboard.org/, and Gerrit.
Figure 1: Code review process in Gerrit
The two communities (i.e., OpenStack and Qt), from which the four open-source software (OSS) projects were selected in this work, use Gerrit as the code review tool. Figure 1 illustrates the modern code review process in Gerrit, which is a well-known code review tool and widely used in modern code review [30]. Gerrit allows for efficient collaboration on code changes before they are merged into the main code base. Once a developer commits new code changes (e.g., software patches) and their descriptions to Gerrit, the system creates a page to record all changes and relevant descriptions (e.g., commit messages). Then, the system conducts a sanity check to verify whether the patch is compliant and to ensure that the code does not contain obvious compilation errors. After the submitted patch passes the sanity check, code reviewers manually examine the patch and provide comments to correct any potential errors, followed by giving a voting score. The comments are subsequently used by the developers to improve the patch and submit it for another review. Finally, the submitted patch is merged into the code repository after passing integration tests, which confirms that there are no issues or conflicts. Alternatively, the patch can be abandoned if the voting is negative.
The above process is used, among others, in revisions that enable the analysis of potentially significant architectural changes. Code review comments provide software developers a finer granularity for investigating architectural changes and violations in their daily development routines [31]. Meanwhile, developers provide reasoning and rationale for the changes they make in code review comments, which enables code review data to be a valuable source of knowledge for explaining changes [32], offering insights into developers’ concerns, suggestions, and potential violations.
3 Study Design
This section describes the research questions that motivated this research, followed by a detailed explanation of the data collection and the ML and DL models to automatically identify architecture violations from textual artifacts. Finally, we provide details of the validation survey and interview conducted to assess the effectiveness of automated identification of violation symptoms in practice.
Figure 2 provides an overview of the whole process. As illustrated in this figure (see the top box, Dataset Creation), we collected data and established a dataset on architecture violations in our previous work [11]. During Phase 1, we explored the possibility of using this dataset to train ML/DL-based classifiers to automatically identify violation symptoms in code review comments. During Phase 2, we sent the survey questionnaire to 200 involved participants and received 24 responses that were used to validate the usefulness of automated identification of violation symptoms in practice.
Figure 2: An overview of the research process
3.1 Research Questions
In this work, we aim to explore the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. Specifically, we formulated the goal of this study by following the Goal-Question-Metric method [33]: analyze code review comments for the purpose of exploring the feasibility of automatically identifying violation symptoms of architecture erosion with respect to the performance on architecture erosion identification from the point of view of developers in the context of open source software development. To achieve our goal, we formulate three Research Questions (RQs):
RQ1: What is the performance of the classifiers in identifying violation symptoms from code review comments? RQ1.1: Which classifier performs best in identifying the violation symptoms? RQ1.2: Which word embedding models can help to improve the performance of classifiers? RQ1.3: To what extent, does the dimensionality of word embedding models impact the performance of classifiers?
Rationale: This RQ focuses on training feasible classifiers to automatically identify violation symptoms from textual artifacts. As mentioned in Section 1, the percentage of violation symptoms (belonging to architecture-level issues) is relatively lower than code-level issues. We thus expect that the description of violation symptoms is scattered in textual artifacts during development. RQ1.1 focuses on finding out the classifier with the best performance to distinguish violation symptoms from natural language by comparing several popular ML and DL algorithms. RQ1.2 aims at employing several well-known word embedding models to generate word vectors as features of labeled data, in order to choose the word embedding models that can help to improve the accuracy of classifiers. The word embedding models are a common ground for both ML and DL-based classifiers in our study. Finally, with RQ1.3, we seek to compare the differences in dimension size among different word embedding models, as it is one of the most important hyperparameters affecting the model performance[34]. Specifically, RQ1.3 endeavors to shed light on the extent to which the selection of dimensionality impacts the performance of the classifiers in automatically identifying violation symptoms.
RQ2: To what extent, a voting strategy can help to improve the performance of identifying violation symptoms from code review comments?
Rationale: After training individual classifiers based on ML and DL algorithms, we want to know whether, and to what extent, an ensemble classifier composed of trained classifiers can help to improve the performance of identifying violation symptoms from code review comments. To answer this RQ, we compared the performance of individual classifiers with that of a combined classifier. More specifically, we employed a voting strategy to integrate the prediction values of each classifier and generated results after voting among individual classifiers. Answering this RQ can help to explore the possibility of further improving the performance of automated identification of violation symptoms.
RQ3: Do practitioners find automated identification of violation symptoms useful in practice?
Rationale: If the results generated by the classifiers based on ML and DL algorithms are acceptable regarding the effectiveness of identified violation symptoms, automatically identifying violation symptoms from textual artifacts would be helpful in practice. Such automated techniques could facilitate tasks such as locating architecture violations and inspiring further investigation and remediation of similar architectural issues. To answer this RQ, we conducted a survey and semi-structured interviews to investigate how practitioners (both developers and reviewers) assess the effectiveness and applicability of the trained classifiers.
3.2 Data Collection
We relied on the dataset that we mined and manually identified in our previous study [11]. As shown in Figure 2 (see Dataset Creation), we followed five steps to establish a specific dataset on architecture violations mined from code review comments. More specifically, we mined code review comments of four OSS projects in the OpenStack (i.e., Nova and Neutron) and Qt (i.e., Qt Base and Qt Creator) communities between 2014 and 2020. We mined the data through the REST API [^4] https://gerrit-review.googlesource.com/Documentation/rest-api.html supported by Gerrit. We first conducted experiment-based keyword refinement to improve our keyword search, as our previous experience indicated that a keyword-based approach could only capture very small proportions of code review comments containing violation symptoms [11]. Subsequently, we proceeded with manual identification, and pilot and formal data labeling of review comments related to architecture violations. In total, we manually identified and labeled 606 code review comments pertaining to architecture violations. Code review comments were created for patches, and comments were made by code reviewers and developers. In terms of the data labeling process, four researchers first conducted a pilot data labeling by randomly selecting 50 review comments, to reach a consensus (the inter-rater agreement between the researchers measured in the Cohen’s Kappa coefficient value [35] was 0.857) and ensure that we have the same understanding of violation symptoms. Any disagreements were discussed between the four researchers to reach a consensus. After that, the four researchers started the formal data labeling by dividing the retrieved 21,583 code review comments into four parts (each researcher manually labeled around 5,400 comments). After the formal data labeling, the first author checked the data labeling results from the other three researchers to make sure that there were no false positives. To mitigate potential bias, the four researchers discussed all the conflicts in the labeling results until we reached an agreement. To summarize, all data labeling results were checked by at least two researchers and conflicts were discussed among four researchers. Subsequently, all the collected code review comments regarding architecture violations (i.e., 606 code review comments) from the four projects were combined into an integrated dataset used in this work. The dataset has been made available in a replication package [36].
3.3 Phase 1 - Automatic Classification of Violation Symptoms
In this section, we describe the experimental setup that we followed to evaluate the performance of the trained classifiers (see Figure 3). The steps in this setup correspond to the steps of Phase 1 in Figure 2. First, we pre-processed the collected code review comments to generate structured word sets, and then utilized word embedding techniques to encode the words and generate vectors for presenting the words in review comments. Subsequently, the generated vectors acted as the input to train the classifiers based on ML/DL models, in order to learn to classify review comments as violations and non-violations. The experimental environment is a computer equipped with Intel(R) Core(TM) i7-10510U CPU and 16GB RAM, running Windows 11 (64-bit). The next subsections elaborate on these steps for both the ML and DL models. More details of the experiments (such as hyperparameters and experimental environment) can be found in the replication package [36].
Figure 3: The framework of the experimental setup for classifying violation symptoms of architecture erosion from code review comments
3.3.1 Machine Learning Models
Before training ML models, we need to pre-process our collected data, which involves transforming raw data into a structural format for model training. The pre-processing includes five steps listed as follows:
Step 1. Tokenization: The process of tokenization is to break a stream of text into words, punctuation, and other meaningful elements called tokens. In this work, we tokenized the review comments by splitting the text into its constituent set of words.
[…]
Step 3. Stop words Removal: Stop words are frequently observed in text, for the purpose of distinguishing between texts, such as “the”, “are”, and “is”, but they are typically devoid of valuable semantic content. Thus, we removed stop words in our dataset by using the Natural Language Toolkit (NLTK) [37].
Step 4. Capitalization Conversion: In order to ensure the consistency of word form and to avoid redundant counts of words, all textual data was converted to lowercase.
Step 5. Stemming: Stemming can be used to convert the words to their base forms. For instance, “architecture” and “architectural” have the same stem (i.e., token) “architectu”. We conducted a stemming process (using Snowball Stemmer from the NLTK toolkit [37]) to obtain the stem of each token, since it is a more comprehensive and aggressive version of the Porter Stemmer [38].
Word Embedding. Feature extraction techniques are widely utilized to transform unstructured text sequences into structured feature spaces (e.g., matrices, vectors, or encodings). In this work, we employ word embedding techniques to extract textual features for generating interpretable input for machine learning algorithms. Word embedding is a prominent method employed in natural language processing to represent words through dense vectors. In this work, we utilized the three prevalent and popular techniques (i.e., word2vec, fastText, and GloVe) to generate word embeddings. These techniques have gained widespread use in machine learning and deep learning-based tasks, owing to their efficacy in capturing the semantic and syntactic relationships among words.
  • word2vec: word2vec was developed in 2013 [39] and it takes words as input and provides a word embedding matrix that contains high-dimensional unique vectors to represent each distinct word in given training corpora. word2vec can generate word embeddings through skip-gram and Continuous Bag-of-Words (CBOW) models. In this study, we adopted the pre-trained word2vec model proposed by Efstathiou et al. [40]. This model was trained based on the corpus comprising over 15 GB of textual data collected from Stack Overflow posts, which contains a plethora of textual information in the software engineering domain.
  • fastText: fastText is an open-source and lightweight library proposed by Meta AI Research lab in 2017 [41] and it is capable of learning text representations and training text classifiers. Based on the fastText model, Meta published pre-trained word vectors with 300 dimensions for 157 languages constructed on Common Crawl and Wikipedia [^5] https://fasttext.cc/docs/en/crawl-vectors.html. In this study, we employed the pre-trained fastText model, which can be transformed into 100 and 200-dimensional models, in order to control variables and avoid the impact of dimensionality on classifiers’ performance.
  • GloVe: Global Vectors for Word Representation (GloVe) is another common word embedding technique [42]. A GloVe model is inspired by aggregated word co-occurrence statistics that may encode global information for words. Compared with word2vec, GloVe avoids the weakness regarding using local word co-occurrence information. In this study, we adopted the GloVe pre-trained word vectors on Twitter data (two billion tweets) with 200 dimensions.
Model Training. Given the binary classification task in this study, we selected five common ML algorithms that are widely used in classification tasks [43], including Support Vector Machine (SVM), Logistic Regression (LR), Decision Tree (DT), Bernoulli Naive Bayes (NB), and k-Nearest Neighbor (kNN). These classification algorithms are also commonly applied in many areas of software engineering studies [44].
For the ML classifiers, we performed a grid search to tune the hyperparameters of each model and picked the best-performing ML models with specific hyperparameters on the validation set. Grid search, an exhaustive search technique, explores every possible combination within a predefined search space to identify the optimal configuration. More specifically, this method is widely employed in the realms of machine learning and optimization and it involves a systematic evaluation that examines all of the combinations of a set of candidate settings to find the best hyperparameter combination [34].
3.3.2 Deep Learning Models
Similarly to the ML models, for the DL models we pre-processed the collected data before training DL classifiers, using the same pre-processing steps described in Section 3.3.1. For the pre-processing of DL classifiers, we needed to ensure the equal size of each paragraph length; therefore, we utilized the zero-padding strategy before dimension transformation to equalize all input text (i.e., the maximum length is set to 2000 words).
Model Training. In recent years, DL-based models are increasingly used in various text classification tasks and Convolutional Neural Network (CNN) has become one of the most popular model architectures for text classification [45]. We selected the TextCNN model to train the DL-based classifiers, since TextCNN, as an adapted CNN-based model, is reported to improve the state of the art on text classification [45]. TextCNN is a CNN-based architecture designed for processing textual data, and was introduced by Yoon Kim in 2014 [46]. It can be used for a variety of Natural Language Processing (NLP) tasks, such as text classification and sentiment analysis. Compared with traditional ML models regarding text processing, the key advantage of TextCNN is its ability to automatically learn features from raw textual data, without relying on manual feature engineering.
We built the TextCNN for identifying violations and non-violations in code review comments through several processing layers, including an input layer, a convolutional layer, a pooling layer, a fully connected layer, and an output layer. Specifically, the TextCNN model applies one layer of convolutions over n-dimensional word embeddings to capture n-gram features at different scales. The convolutional layer learns to detect local patterns in the input text. The resulting feature maps are then fed through one max-pooling layer, which extracts the most salient features from each map and concatenates them to produce a fixed-length vector representation of the input text. This vector is then passed through one or more fully connected layers to produce the final output, i.e., the predicted labels for the input review comments.
TextCNN is a convolutional neural network model that is commonly used for text classification tasks. This model takes a sentence as input and applies multiple convolutional filters to extract important features from the text. A TextCNN classifier based on the vocabulary set of a dataset uses a one-hot encoding scheme to represent words and generate the word embedding, while a TextCNN classifier can also utilize pre-trained word embedding models to capture semantic information of words. Additionally, the TextCNN model requires a proper setting to construct neural network models, such as hyperparameters and word embedding parameters. In this work, we set the embedding dimension to 200, the learning rate to 0.001, and the dropout to 0.25 to handle the overfitting, while the filter region size is (3, 4, 5). Each DL model’s batch size is set to 16, and the network is trained for 100 epochs with the early stopping strategy with the purpose of avoiding overfitting. We configured the patience parameter to 8, indicating that the training process would terminate after 8 consecutive epochs with no observed improvement in model performance. These approaches are common practices in machine learning, aimed at preventing overfitting and optimizing the model’s convergence [47].
3.3.3 Performance Evaluation
For the performance evaluation of ML/DL classifiers, we adopted four commonly used metrics, namely, Precision, Recall, F1-score (i.e., the harmonic mean of Precision and Recall), and Accuracy.
Precision presents the proportion of correctly classified violation symptom comments to all comments classified as violation symptoms. Recall represents the proportion of all violation symptom comments that are correctly identified. F1-score is the harmonic mean between precision and recall, assessing if an increase in one compensates for a decrease in the other. Accuracy is the percentage of correct classifications by a trained learning model. The above metrics are defined and calculated by the following equations.
$\rm Precision=\frac{TP}{TP+FP}$ (1)
$\rm Recall=\frac{TP}{TP+FN}$ (2)
$\rm F_{1}-score=2\times\frac{Precision\times Recall}{Presion+Recall}$ (3)
$\rm Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (4)
Here, we need to use four statistics to calculate the metrics, namely, False Positives (FP) represent the number of non-violation comments that are classified as violation comments; False Negatives (FN) represent the number of violation comments that are classified as non-violation comments; True Positives (TP) represent the number of violation comment that are classified as violation comments; True Negatives (TN) represent the number of non-violation comments that are classified as non-violation comments.
To avoid the class imbalance problem [48], a data distribution problem that significantly impacts classifiers’ performance, we constructed our dataset with a balanced class distribution (i.e., the same number of review comments labeled as “violations” and “non-violations”), which avoids the poor generalization ability of classifiers. With respect to the model training and validation strategy, we employed 10-fold cross-validation when we conducted a performance evaluation of the classifiers. We divided the collected dataset into three parts, i.e., we randomly set aside 80% of the data for the training set (60%) and validation set (20%), and 20% for the test set. Besides, we set a fixed seed for reproducibility during experiments.
3.4 Ensemble Classifier
To identify architecture violations from textual artifacts (i.e., code review comments), we trained several ML and DL-based classifiers to predict the labels of new review comments. However, the trained classifiers have varying performances in predicting the labels of target artifacts. In RQ2, we asked whether composing the individual classifiers into an ensemble classifier can help improve the overall performance. This composition can be done by utilizing the ensemble learning technique, i,e., majority voting strategy [49]. Specifically, a combined classifier of several individual classifiers is generated with the purpose of building an improved classifier, which predicts an output (class) based on the highest probability of chosen class as the output [49]. This aggregating criterion is like an election process and can produce combined voting results from each classifier’s output.
The voting criteria include Hard Voting (voting is calculated on the predicted output class) and Soft Voting (voting is calculated on the predicted probability of the output class). To reduce the training time, we employed the Hard Voting (a simple majority voting strategy) strategy in this work. In other words, the final predicted results depend on the majority output (i.e., labels) of the individual classifiers.
3.5 Phase 2 - Validation: Survey and Interview
To answer RQ3, we designed and executed a survey and semi-structured interviews by following the guidelines proposed by Kitchenham and Pfleeger’s personal opinion surveys [50]; in terms of the survey, we used an anonymous, web-based questionnaire to increase the response rate.
(1) Survey. Surveys can be conducted in different ways, such as online questionnaires, phone surveys [51]. Our survey is an online questionnaire of the cross-sectional type, which allows us to collect data and compare differences between participants at a single point in time. We adopted self-administered questionnaires to collect responses, since (a) respondents can answer the survey at their convenience (reducing our time investment and increasing the response rate); (b) questionnaires can reduce research bias since the researcher does not have any contact with the respondents when they fill out the survey; and (c) respondents enjoy better privacy than interviews.
We invited participants through customized emails (see the replication package [36]) attaching the links to their code review comments that contain violation symptoms and are identified by our approach, as well as the survey form (also see the replication package [36]). Our survey was constructed based on four parts: Problem Statement, Our approach, Survey Goal, and Survey Questions. Problem Statement introduces the research background regarding what is architecture erosion and its violation symptoms, as well as why it is important to investigate violation symptoms. Our approach describes the approach that we developed to automatically identify violation symptoms with high performance from a large number of code reviews as well as two examples of violation symptoms of architecture erosion detected by our approach from code review comments. In Survey Goal, we explain the purpose of this survey. In Survey Questions, we ask five general questions and present six statements concerning the usefulness of our approach. Note that, before the formal survey, we did a pilot survey (receiving 6 responses) with the purpose of revising and improving our survey questions and examples.
The five general survey questions include demographic information (e.g., their countries and work experience). The six statements are assessed using a five-point Likert scale (Strong Agree, Agree, Neutral, Disagree, and Strongly Disagree) with one option (“I don’t know”). This scale enables participants to rate the extent that they agree or disagree with the statements. Finally, at the end of the survey, we ask an open-ended question to let the respondents freely provide comments and suggestions about our survey. To formulate the statements, we took inspiration from the survey questions of two survey studies [52, 53] that evaluate the usefulness of automatic classifiers based on ML and DL techniques in the software engineering community. We chose to use declarative statements in our validation survey since declarative statements are deemed appropriate for exploratory studies to convey simple and clear judgements; in contrast, descriptive statements are more suitable in explanatory studies to provide the details of the judgements in various aspects. The statements are as follows:
  • Statement 1. “Violation symptoms in code review comments identified by the approach, represent potential architectural violations”.
  • Statement 2. “Violation symptoms in code review comments identified by the approach can be used to improve the code quality”.
  • Statement 3. “I (as a practitioner) can find potentially constructive and useful architectural information from violation symptoms in code review comments identified by the approach”.
  • Statement 4. “Violation symptoms in code review comments identified by the approach can help us identify violation-related issues faster than if we do this manually”.
  • Statement 5. “Violation symptoms in code review comments identified by the approach might help us locate, identify, and prioritize potential architectural violations in our systems”.
  • Statement 6. “Violation symptoms in code review comments identified by the approach can provide us with input to find other violations of similar nature or otherwise”.
(2) Interview. As a supplement to the survey, we also invited the same practitioners to participate in semi-structured interviews, which allow for more in-depth discussions. Specifically, the interviews focused on the developers’ opinions on the effectiveness of our trained classifiers, as well as the role of such automated techniques in the code review process during software development. The semi-structured interviews were conducted by focusing on the following three open-ended Interview Questions (IQs):
  • IQ1: How do you think the integration of an automated system for identifying violation symptoms of architecture erosion from code review comments would fit within the code review workflow? This system would analyze reviewer comments in real-time to flag potential or overlooked violation symptoms. By doing so, it would prevent the incorporation of code changes that violate system architecture, thereby contributing to a more robust and reliable development process. What are your thoughts on the potential benefits and challenges of such integration?
  • IQ2: Based on your experience, what strategies would you recommend to effectively use automated identification of violation symptoms in the code review process? How can this technology be utilized to maximize its benefits and efficiency in identifying and addressing architectural issues during code review?
  • IQ3: In addition to code review scenarios, can you envision other contexts or scenarios where our approach for automated identification of violation symptoms could be beneficial in software development practice? What potential applications or integrations do you see as valuable in enhancing code quality and development workflows?
Participants. Our collected data from Gerrit contains the developers’ information (i.e., names and email addresses); this information is not included in the replication package for privacy reasons. For the survey, we invited participants to fill out our survey by sending emails to the developers whose comments are identified by our classifiers. During the pilot survey, we got responses from 6 developers. The pilot survey helped us to evaluate and refine our survey design. For example, we added more precise examples of violation symptoms from code review comments in our survey in order to enhance practitioners’ understanding, and we refined the content of the customized emails based on the received feedback from the pilot survey participants. Then, we invited the remaining developers involved in the discussions on violation symptoms to fill in the survey. Finally, for the semi-structured interviews, we invited the above-mentioned participants to give feedback on our trained classifiers and the role of such automated techniques during code review. We first conducted two pilot interviews and found that, the interview questions are suitable. Subsequently, we interviewed the remaining participants who accepted our invitation, resulting in a total of six interviews.
Sample and Population. The target population is limited to developers whose code review comments are identified by our approach. We used a Non-Probabilistic Sampling method in this survey, namely, Convenience Sampling. As a type of non-probability sampling method, Convenience Sampling can help us to conveniently reach a suitable number of developers who are willing to participate in our survey. The reason we chose this sampling method is that we could not intentionally choose a sample due to the limitations of geographical constraints and scheduling conflicts.
Data Analysis. For the survey questions, we applied descriptive statistics to analyze the responses to the demographic and Likert scale questions. For the open-ended interview questions, interviews were recorded and transcribed; subsequently, we used the open coding and constant comparison method [54] to analyze and categorize the qualitative responses.
4 Results
4.1 RQ1: Identifying Violation Symptoms
To answer RQ1, we trained five ML classifiers and one DL classifier based on three word embedding techniques (i.e., word2vec, fastText, and GloVe). Moreover, we evaluated the performance of these classifiers and compared the difference between their performance.
TABLE II: Performance comparison among ML-based classifiers
Classifier Precision Recall F1-score Accuracy
w2v FT GloVe Average w2v FT GloVe Average w2v FT GloVe Average w2v FT GloVe Average
SVM 0.789 0.728 0.759 0.759 0.828 0.746 0.828 0.801 0.808 0.737 0.792 0.779 0.803 0.734 0.783 0.773
LR 0.832 0.777 0.798 0.802 0.730 0.713 0.713 0.719 0.777 0.744 0.753 0.758 0.791 0.754 0.766 0.770
NB 0.728 0.743 0.717 0.729 0.615 0.615 0.623 0.618 0.667 0.673 0.667 0.669 0.693 0.701 0.689 0.694
DT 0.667 0.657 0.643 0.656 0.689 0.721 0.664 0.691 0.677 0.688 0.653 0.673 0.672 0.672 0.648 0.664
kNN 0.863 0.788 0.797 0.816 0.516 0.549 0.451 0.505 0.646 0.647 0.576 0.623 0.717 0.701 0.668 0.695
Average 0.776 0.739 0.743 0.753 0.676 0.669 0.656 0.667 0.715 0.698 0.688 0.700 0.735 0.712 0.711 0.719
  • w2v: word2vec, FT: fastText
TABLE III: Performance comparison among DL-based classifiers
Classifier Precision Recall F1-score Accuracy
TextCNN_voc 0.475 0.541 0.469 0.471
TextCNN_w2v 0.584 0.680 0.596 0.598
TextCNN_FT 0.728 0.549 0.667 0.672
TextCNN_GloVe 0.507 0.885 0.434 0.512
Average 0.574 0.664 0.542 0.563
  • voc: vocabulary set, w2v: word2vec, FT: fastText
TABLE IV: Performance comparison of the classifiers based on the fastText model with different dimension size
Classifier Precision Recall F1-score Accuracy
100-dim 200-dim 300-dim 100-dim 200-dim 300-dim 100-dim 200-dim 300-dim 100-dim 200-dim 300-dim
SVM 0.745 0.728 0.657 0.672 0.746 0.754 0.707 0.737 0.702 0.721 0.734 0.680
LR 0.704 0.777 0.745 0.664 0.713 0.648 0.684 0.744 0.693 0.693 0.754 0.713
NB 0.745 0.743 0.712 0.623 0.615 0.607 0.679 0.673 0.655 0.705 0.701 0.680
DT 0.641 0.657 0.612 0.615 0.721 0.648 0.628 0.688 0.629 0.635 0.672 0.619
kNN 0.831 0.788 0.831 0.525 0.549 0.443 0.643 0.647 0.578 0.709 0.701 0.676
TextCNN 0.577 0.728 0.494 0.648 0.549 0.730 0.585 0.667 0.461 0.586 0.672 0.492
TABLE V: Performance comparison of ensemble classifiers
Classifier Precision Recall F1-score Accuracy
Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B
SVM 0.759 0.789 0.795 4.74% 0.76% 0.801 0.828 0.828 3.37% 0.00% 0.779 0.808 0.811 4.11% 0.37% 0.773 0.803 0.807 4.40% 0.50%
LR 0.802 0.832 0.833 3.87% 0.12% 0.719 0.730 0.738 2.64% 1.10% 0.758 0.777 0.783 3.30% 0.77% 0.770 0.791 0.795 3.25% 0.51%
NB 0.729 0.743 0.738 1.23% -0.67% 0.618 0.615 0.623 0.81% 1.30% 0.669 0.673 0.676 1.05% 0.45% 0.694 0.701 0.701 1.01% 0.00%
DT 0.656 0.657 0.689 5.03% 4.87% 0.691 0.721 0.746 7.96% 3.47% 0.673 0.688 0.717 6.54% 4.22% 0.664 0.672 0.705 6.17% 4.91%
kNN 0.816 0.863 0.836 2.45% -3.13% 0.505 0.516 0.500 -0.99% -3.10% 0.623 0.646 0.626 0.48% -3.10% 0.695 0.717 0.701 0.86% -2.23%
TextCNN 0.574 0.728 0.537 -6.45% -26.24% 0.664 0.549 0.713 7.38% 29.87% 0.542 0.667 0.613 13.10% -8.10% 0.563 0.672 0.549 -2.49% -18.30%
TABLE VI: Performance comparison among ensemble ML-based classifiers with three word embeddings
Classifier Precision Recall F1-score Accuracy
Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B
ML_word2vec 0.776 0.789 0.810 4.38% 2.66% 0.676 0.828 0.697 3.11% -15.82% 0.715 0.808 0.749 4.76% -7.30% 0.735 0.803 0.766 4.22% -4.61%
ML_fastText 0.739 0.777 0.798 7.98% 2.70% 0.669 0.713 0.713 6.58% 0.00% 0.698 0.744 0.753 7.88% 1.21% 0.712 0.754 0.766 7.58% 1.59%
ML_GloVe 0.743 0.759 0.768 3.36% 1.19% 0.656 0.828 0.705 7.47% -14.86% 0.688 0.792 0.735 6.83% -7.20% 0.711 0.783 0.746 4.92% -4.73%
RQ1.1: Performance of Classifiers
For ML-based classifiers, as shown in Table II, we trained the five types of ML classifiers with three word embedding techniques (i.e., word2vec, fastText, and GloVe), leading to 15 classifiers (i.e., the five ML models times three word embeddings). We calculated the average values of the classifiers over the three pre-trained word embeddings. For each classifier based on three word embeddings, we underlined the best result of each metric (horizontal comparison); for each metric of the classifiers (precision, recall, F1, and accuracy), we marked the best metric result of each classifier on three word embeddings in bold (vertical comparison). From the results, we can see that the kNN classifiers have better scores of Precision than other classifiers with an average Precision of 0.816. As for the remaining metrics, the SVM classifiers achieve relatively better performance on nearly all metrics than other classifiers, with an average Recall of 0.801, F1-score of 0.779, and Accuracy of 0.773. Specifically, the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803.
For DL-based classifiers, as shown in Table III, we trained one type of classifier, i.e., TextCNN, based on the vocabulary set from our dataset and three word embedding techniques (i.e. in total four classifiers) as mentioned in Section 3.3.1. We can see that TextCNN_FT performs the best compared with the other three DL models.
The ML-based classifiers demonstrated superior performance, achieving, on average, a Precision of 0.753, a Recall of 0.667, an F1-score of 0.700, and an Accuracy of 0.719. In contrast, the DL-based classifiers exhibited lower performance, with an average Precision of 0.574, Recall of 0.664, F1-score of 0.542, and Accuracy of 0.563. These findings suggest that the ML-based classifiers are more effective in accurately classifying the target variable than the DL-based classifiers. Overall, among these classifiers, the SVM classifiers outperform other classifiers, exhibiting the best average results on all metrics, except for the Precision metric.
RQ1.2: Comparison of Embedding Models
From the results in Tables II and III, we can also observe the differences between the classifiers based on three pre-trained word embedding models. As mentioned before, for each classifier based on three word embeddings, we underlined the best result of each classifier on each metric (horizontal comparison). When using the word2vec pre-trained word embedding, most of the ML classifiers have relatively better performance on Precision and Accuracy. Most of the ML classifiers with the fastText model have fairly better F1-scores. Besides, we found that the TextCNN classifier based on the fastText pre-trained word embedding achieves the best performance on all metrics except for Recall, with a Precision of 0.728, a Recall of 0.549, an F1-score of 0.667, and an Accuracy of 0.672.
In addition, we can see in Table III that the overall performance of the three TextCNN classifiers with pre-trained word embeddings outperforms the TextCNN classifier based on the vocabulary set from our dataset. This is because pre-trained word embedding models contain more features due to the large textual data used for generating word vectors, and they have been proven to be invaluable for improving the performance in NLP tasks [55, 56].
In general, the three pre-trained word embedding models have an impact on the performance of the DL-based classifiers. As shown in Table III, when we compared their classification performance, in most cases, the classifiers with the fastText model can achieve relatively good performance.
RQ1.3: Comparison of Word Embedding Model Dimensionality
According to the results of RQ1.2, the fastText model can help to generate better results than the other two pre-trained word embedding models. Therefore, to further explore the impact of dimension sizes of the pre-trained word embedding models, we trained the classifiers based on the fastText model and chose three common dimension sizes: 100, 200, and 300. In Table IV, similarly as before, we underlined the best result of each classifier on each metric (horizontal comparison). We can see that five of the six ML/DL classifiers can generate better F1-scores based on the fastText model with 200-dimensional word embedding, compared with 100 and 300-dimensional models.
RQ1 Summary. The SVM classifier trained on the pre-trained word2vec word embedding can achieve the best performance on identifying violation symptoms, with an F1-score of 0.779. In most cases, the classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200 as the dimension size of the pre-trained word embedding models can generate better performance than 100 and 300-dimensional models.
4.2 RQ2: Improving Performance with Voting
To answer RQ2, we employed an ensemble learning technique, i.e., majority voting strategy, to build an ensemble classifier. We compared the performance of the individual classifiers with the ensemble classifiers. More specifically, we explored to what extent combining one type of classifier based on three word embedding techniques can achieve better performance, as shown in Table V and Table VI.
In Table V, the “Mean” column presents the Average values of each metric of each classifier on three word embeddings from Table II and Table III. The “Best” column denotes the corresponding classifier with the best performance from Table II and Table III. The “Voting” column refers to the performance of the ensemble classifier (based on classifiers with three word embeddings) on each metric after utilizing a majority voting strategy. For example, the Voting values of the SVM classifiers in Table V refer to the voting results among the three individual classifiers SVM_w2v, SVM_FT, and SVM_GloVe. “Impro_M” and “Impro_B” represent the improved results of the corresponding ensemble classifiers compared with the results in “Mean” and “Best” columns, respectively.
According to Table V, when we utilized the voting strategy for each ML classifier based on the three word embeddings (i.e., word2vec, fastText, GloVe), we can see that all ensemble classifiers perform better than the mean values of individual classifiers, with F1-score improvement from 0.48% to 13.10%. Even compared with the best-performing classifiers, most of the ensemble classifiers outperform the individual ML/DL classifiers, except for kNN. Moreover, the ensemble TextCNN of the three TextCNN classifiers based on pre-trained word embeddings (see Table III) does not perform as well as TextCNN_FT (i.e., the best one) on Precision, F1-score, and Accuracy. This means that TextCNN_FT with the fastText pre-trained word embedding performs best compared with the other three TextCNN classifiers, and the voting strategy does not have an obvious improvement for the trained DL classifiers.
In addition, as shown in Table VI, we utilized the voting strategy for the five ML classifiers based on each of the three word embeddings. For example, ML_word2vec is the ensemble classifier of the five ML classifiers based on the word2vec word embedding (i.e., SVM_w2v, LR_w2v, NB_w2v, DT_w2v, and kNN_w2v). Similarly as before, the “Mean” column refers to the average metric results of ML classifiers (see the last row in Table II); the “Best” column presents the best-performing classifiers in Table II; the “Voting” column denotes the performance results of the ensemble classifiers after voting. It is clear that the three ensemble classifiers outperform the average performance results of the individual classifiers, with improvements from 3.11% to 7.98%. Even compared with the best-performing individual classifiers (i.e., the maximum values of the voting samples), the ensemble classifier ML_fastText can still demonstrate performance improvement on all metrics.
Moreover, we can observe the performance difference of ensemble classifiers with three word embeddings. In general, for the selected five ML models, the fastText pre-trained word embedding model can achieve relatively better performance improvements on almost all the metrics compared with the models using word2vec and GloVe. This finding affirms a similar observation in the result of RQ1.2.
RQ2 Summary. In comparison to individual ML classifiers, the ensemble classifiers achieve better performances after utilizing the majority voting strategy. However, no significant improvement was observed in the performance of the DL classifiers on our dataset when implementing the ensemble technique.
4.3 RQ3: Validation in Practice
As elaborated in Section 3.5, we designed and conducted a survey and semi-structured interviews to solicit the perceptions from the practitioners on the usefulness of our trained models to automatically identify violation symptoms from code review comments.
Overall, we planned to send customized survey invitation emails to 200 code reviewers who were involved in the discussion of the collected review comments concerning architecture violations. However, we could only derive valid email addresses for 169 of them. We finally obtained 24 valid responses to our survey and interviewed 6 participants after sending 169 survey invitation emails and reminder emails; the survey response rate of 14.2%, is a reasonable response rate considering that the general response rate is around 5% in empirical software engineering research [57]. Table VII shows the demographic information of the 24 respondents. Twenty two of the respondents have more than 10 years of experience in software development, while two of them have 6-10 years of experience. Among the respondents, 50.0% (12) are software engineers, 33.3% (8) are developers, and 16.7% (4) are architects. Note that, software engineers usually have a broader work range than developers; for example, the former may work on full-stack development while the latter could deal only with the back end. Regarding the size of their companies, ten of the respondents reported that their companies have more than 1000 employees, five respondents worked in companies with employees between 501 and 1000, six respondents in companies with employees between 101 and 500, and three respondents in companies with employees between 21 and 100. Besides, the respondents worked in a variety of application domains.
Figure 4 presents the feedback from the 24 respondents on the usefulness of the trained models, i.e., the automated identification of violation symptoms from the textual content of code review comments. According to the results, most of the respondents (66.7%) agreed that the violation symptoms identified by our models represent potential architectural violations (Statement 1). 37.5% respondents agreed that the identified violation symptoms can be used to improve the code quality (Statement 2), while 41.7% respondents had a neutral attitude about this statement. 54.2% of the respondents agreed that the identified violation symptoms contain potentially constructive and useful architectural information (Statement 3).
Statements 3 to 6 investigate whether the trained classifiers can play an auxiliary role in development and code reviews. Approximately 41.7% respondents strongly agreed (12.5%) or agreed (29.2%) that the identified violation symptoms can help identify violation-related issues faster than manual identification (Statement 4); this indicates that the classifiers may help to speed up code review. Moreover, the majority (62.5%) of the respondents strongly agreed (4.2%) or agreed (58.3%) that automated identification can help locate and prioritize architectural violation issues (Statement 5). Furthermore, 41.7% of the respondents agreed that the identified violation symptoms can inspire them to find out other (or similar) violation symptoms (Statement 6). Notably, one respondent strongly disagreed with all six statements, indicating that he/she might not believe that ML and DL-based classifiers can be useful for identifying architectural violations in practice; we did not treat this response as an outlier. We further analyzed the responses for the six statements and found no significant statistical differences between the identified violation symptoms. Overall, Figure 4 shows that positive feedback (strongly agree + agree) surpasses negative (disagree + strongly disagree) and neural feedback in most of the statements, except for Statement 2.
In terms of the semi-structured interviews, the six practitioners that accepted our invitation are denoted as I1 to I6 in Table 7. Each interview lasted approximately 30 to 40 minutes, resulting in a total of 189 minutes of audio recordings. Furthermore, the transcripts of these interviews spanned 35 pages with over 22K words. In general, all of them expressed a positive attitude towards such an automated way to identify violation symptoms from code review comments. We present the key common points of the interviewees’ opinions below:
  • (IQ1) Potential benefits: The integration of such an automated identification of violation symptoms of architecture erosion could help to reduce the time and effort of developers, which is recognized by four interviewees. As I3 stated, “extra checking could make our work faster and facilitate the code review workflow”. Additionally, I4 mentioned that “Code reviewers prioritize two main categories of issues: those related to functionality and those concerning the architecture, and such an approach is particularly valuable and beneficial for identifying and addressing architectural problems”.
  • (IQ1) Challenges: Three out of six interviewees raised concerns regarding data leakage problems when employing or integrating such an automated approach supported by third-party libraries or specific Large Language Models (LLMs), especially for inner-source projects. Besides, two out of six interviewees expressed their concerns about false positives; as I1 mentioned: “Any tool that gives you warnings that are false, after a while you start ignoring its warnings”. Regardless of the type of automated techniques employed, rich and high-quality training data on violation symptoms and suitable data processing are necessary to ensure the performance and generality of these approaches. As I2 stated, “Training data is quite important for the accuracy of the results, …, there were always garbage outputs even if the input per se is good”. Therefore, putting more effort into improving data processing could pose another challenge for such automated approaches.
  • (IQ2) Strategies: I4 thought that the best practices and specific rules for identifying typical violation symptoms (e.g., layering violations) should be considered and integrated when employing automated techniques during code review. This could help customize such an approach of identifying violation symptoms and then maximize the identification accuracy for inner-source projects inside companies.
  • (IQ3) Visions for improving code quality: With the advent of LLMs, five out of six interviewees are willing to utilize our approach with customized LLMs for specific tasks in the future, e.g., identifying architecture issues by feeding architecture constraints to the LLMs and providing architecture refactoring suggestions.
RQ3 Summary. Practitioners’ responses indicate the usefulness of the automated identification of violation symptoms discussed in code review comments in practice. Besides, the identified violation symptoms can enable and inspire practitioners to find architecturally related violations and prioritize violation-related issues.
TABLE VII: Demographic information of the survey participants and interviewees
# Country Experience Role Company size Company domain
P1 Ireland More than 10 years Software Engineer employees $\geq$ 1000 HCM services
P2 Germany More than 10 years Developer 21 $\leq$ employees $\leq$ 100 Industrial electronics
P3 Finland More than 10 years Software Engineer 21 $\leq$ employees $\leq$ 100 Medical software
P4 Norway More than 10 years Architect 501 $\leq$ employees $\leq$ 1000 Industrial DataOps
P5 Israel More than 10 years Software Engineer employees $\geq$ 1000 Network
P6 Germany 6 - 10 years Developer 101 $\leq$ employees $\leq$ 500 Development tools
P7 USA More than 10 years Developer 501 $\leq$ employees $\leq$ 1000 Iaas
P8 Norway More than 10 years Developer 501 $\leq$ employees $\leq$ 1000 Cross-platform libraries for application development
P9 Norway More than 10 years Developer 101 $\leq$ employees $\leq$ 500 SW development frameworks
P10 USA More than 10 years Software Engineer employees $\geq$ 1000 Software development
P11 Germany More than 10 years Developer 101 $\leq$ employees $\leq$ 500 Middleware
P12 USA More than 10 years Software Engineer employees $\geq$ 1000 Research
P13 (I1) Norway More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Development framework
P14 Finland More than 10 years Architect 21 $\leq$ employees $\leq$ 100 Medical devices
P15 Germany More than 10 years Developer 501 $\leq$ employees $\leq$ 1000 Tools and frameworks for software development
P16 (I2) Germany More than 10 years Developer 101 $\leq$ employees $\leq$ 500 Middleware Platform
P17 Germany More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Defence development
P18 Denmark 6 - 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Network
P19 (I3) Canada More than 10 years Software Engineer employees $\geq$ 1000 Framework tool
P20 India More than 10 years Architect employees $\geq$ 1000 Development framework
P21 (I4) China More than 10 years Architect employees $\geq$ 1000 Cloud, Data, GenAI
P22 USA More than 10 years Software Engineer employees $\geq$ 1000 Tools for software development
P23 (I5) Germany More than 10 years Software Engineer employees $\geq$ 1000 Game development
P24 (I6) Switzerland More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 AI-related research and development
  • P = Survey Participant, I = Interviewee
Figure 4: Practitioners’ responses to statements regarding the usefulness of the trained models
5 Discussion
5.1 Interpretation of Results
5.1.1 RQ1: Identifying Violation Symptoms
Our trained ML/DL classifiers can identify violation symptoms in code review comments. The results of RQ1.1 indicate that certain factors such as the type of classifier, the pre-trained word embeddings, and hyperparameters influence the performance of classifiers. Specifically, we found that the SVM classifier with word2vec outperforms other ML and DL-based classifiers. We also found that DL classifiers do not always have better performance than ML classifiers; we note that training DL classifiers takes much longer time than training ML classifiers, as DL classifiers have much more complex network architecture and more parameters. In recent years, some studies that leveraged DL-based techniques for software engineering tasks have achieved better performance that all other state-of-the-art techniques (e.g., ML-based techniques) [44]. However, the achieved results imply that ML classifiers might be enough for certain text classification tasks with small-size datasets, especially as the trained DL classifiers do not show performance improvement in such a scenario. This finding is also observed in previous studies [58, 59]: for certain software engineering tasks, simple ML-based approaches are capable of achieving equal or even better performance than DL-based approaches within less time. Thus, we advise researchers to not blindly apply DL-based techniques and consider simple approaches first.
Finding 1: ML-based approaches have sufficient performance for many software engineering tasks, such as identification of architecture violations in code reviews.
From the results of RQ1.2, we can see that the TextCNN classifier based on the fastText pre-trained word embedding can perform better than other DL models. We thus emphasize that the fastText pre-trained technique has a better capacity for training well-performed DL models; this also aligns with the finding in the work of Sesari et al. [60], that is, the fastText pre-trained word embedding model has less bias compared to word2vec and GloVe.
Besides, the results of RQ1.3 show that 200-dimensional pre-trained word embedding models outperform 100 and 300-dimensional models. Our results are slightly different from previous studies (e.g., [61, 62]) where the 300-dimensional word embedding model has better performance. One possible reason is that the previous studies only measured the average results of classifiers with different dimensional word embedding models, and the average results might be influenced by the individual results. Our results present and compare the classification performance of each ML/DL classifier, which could be more accurate.
Finding 2: The 200-dimensional fastText pre-trained word embedding model can be used to train classifiers with better performance when conducting binary classification of violation symptoms from code review comments.
In general, due to a large number of hyperparameters of DL classifiers, we did not specifically focus on the best combination of hyperparameters of the TextCNN classifier, but certain key hyperparameters such as the dimension size of word embedding model. Thus, future work can consider performing hyperparameter tuning (e.g., utilizing the grid search technique) to explore other possible combinations of the hyperparameters to further improve performance.
5.1.2 RQ2: Improving Performance with Voting
Evaluating the performance of an ensemble classifier offers insight into the performance improvement compared with individual classifiers. For voting-based approaches, an accurate output is attained only when a majority of the constituent models generate the correct output [63]. Our results provide further empirical evidence for the efficacy of the majority voting strategy regarding improving the performance of text classification. This technique effectively combines the predictions of multiple classifiers to arrive at a final decision, resulting in improved overall classification performance. Nevertheless, we also observed that such performance improvement is not obvious in TextCNN in our case. One possible reason is that we only get an ensemble DL classifier based on three individual TextCNN models, and that causes the voting results to be particularly susceptible to the influence of single results. We argue that the voting strategy remains valuable to minimize potential prediction errors if there are more outputs for voting.
Finding 3: Compared with individual classifiers for identifying architecture violations from code review comments, ensemble learning is an effective way for improving classification performance. In most cases, ensemble classifiers that utilize the voting strategy can enhance all classification performance metrics.
5.1.3 RQ3: Validation in Practice
The existence of architectural violations can be challenging, as they can impede practitioners’ understanding of architecture during the development process. This particularly influences novice developers who lack practical familiarity with architecture. The results of RQ3 indicate a predominantly positive feedback, confirming the usefulness of our trained ML/DL classifiers in potentially helping developers during maintenance and evolution. Through our analysis of the received responses to the statements (e.g., Statements 3 and 5), we found that developers expressed a significant interest in utilizing automated checking approaches, tools, or plug-ins to identify violation symptoms during code review, recognizing their potential to enhance the productivity of dealing with architectural violations through code review. The survey and interview results indicate that developers hold positive attitudes towards such kind of tools. Moreover, such automated techniques regarding identifying potential architecture violations could be supplementary means to minimize architecture violations and improve software quality during code review. For example, such techniques could provide code reviewers with warnings that architecture violations need to be refactored, based on the review discussions. Furthermore, they could help code reviewers prioritize such refactorings based on various criteria (e.g., the importance of involved components, the number of violation symptoms in those components, the seniority of code reviewers). Such a prioritization could help to fix violation symptoms that might otherwise have been ignored.
Moreover, according to the feedback from respondents, certain practitioners might hold a conservative attitude toward ML and DL-based classifiers for identifying violation symptoms, which could be observed from the “Neutral” results in Figure 4, especially for Statement 2. One possible reason could be that such identification techniques merely provide warnings about potential issues, rather than offering refactoring suggestions to effectively address those violation symptoms. Another possible reason is that they might be skeptical about the performance of the classifiers (false positives or negatives), casting doubt about the reliability of identified architectural violations. Nevertheless, no classifier can ever achieve 100 % accuracy in classification tasks, and there is always a trade-off between precision and recall. This disclaimer should be made clear when showing classification tasks to practitioners. Another potential reason could be that certain identified review comments may be part of the architectural design discussions during the development process, rather than the final architectural decisions. Such concerns are beyond the target of our research, as the trained classifiers can only provide judgments based on the collected violations, instead of exploring the evolution of architecture violations.
Finding 4: Although there are concerns about the correctness of the identified architecture violations, the positive responses from participants suggest that practitioners are receptive to the use of automated techniques for identifying architecture violations during code review.
5.2 Implications
Establish and enrich datasets. Many studies utilized or proposed automated approaches to identify or predict knowledge-related information from textual artifacts. However, no studies focused on the automated identification of violation symptoms from code review comments. This suggests that there is a need for more empirical reports and follow-up tools regarding such issues. Our study provides a starting point for researchers to automatically identify architecture violations in textual artifacts such as review comments. On top of this, we encourage researchers and practitioners to establish and share datasets related to architecture violations to lay a solid foundation for future research.
Moreover, our study is scalable in practice, as practitioners can retrain superior classifiers by adding more code review comments of violation symptoms from different projects, either manually or by using our classifiers. Besides, mining architecture violations from multiple textual artifacts (e.g., issues, source code comments, and pull requests) can provide a more comprehensive examination of architectural violations and also enrich the dataset used for identifying such violations. Furthermore, our findings regarding the experiments in this study might benefit or be adapted by researchers and practitioners to identify other issues, such as security and vulnerability discussions.
Suggestion 1: Researchers and practitioners are encouraged to establish and share datasets related to architecture violations. Besides, our study is scalable to retrain superior classifiers by adding more data from different sources, and can also be adapted to identify other issues.
Choose appropriate classifiers. An ensemble classifier is a collection of classifiers that work together to classify instances by combining their individual outputs through voting. Researchers can consider exploring different selections of classifiers to further enhance the performance of ensemble classifiers. For example, to some extent, assigning varying weights to individual classifiers might partially mitigate the impact of biased classifiers on voting results.
Suggestion 2: It is recommended for researchers to explore various combinations of classifiers and assign different weights to individual classifiers to further enhance the classification performance of ensemble classifiers for identifying architecture violations.
Take time investment into consideration. Even though our study did not specifically investigate the time efficiency of classifiers’ training, we observed that the TextCNN classifiers exhibited a long training time due to a large number of parameters. In comparison to traditional ML classifiers, DL classifiers require a longer training time, yet the results did not demonstrate a significant performance improvement. Therefore, when considering the practical implementation of classifiers, it is important to take into account the cost-benefit ratio of “Performance/Time”. In situations where there is a limited dataset, traditional machine learning techniques may suffice in achieving the desired goals. Further research may explore additional methods for balancing the trade-off between performance and training time for classifiers.
Suggestion 3: For certain software engineering tasks, constructing simple classifiers (e.g., individual or ensemble classifiers) can yield a superior return on investment compared to training complex DL-based classifiers, particularly when time investment needs to be taken into account.
Optimize code review practice. In the future, researchers and practitioners can explore the development of plug-ins for code review platforms (e.g., Gerrit) that provide warnings of violations to developers or even suggestions to fix those violations. Such plug-ins can be developed utilizing historical code review data or harnessing the existing large language models through fine-tuning. These plug-ins have the potential to assist developers and maintainers in conducting comprehensive examinations of preexisting architectural violations, and may even serve as reminders of their presence. Additionally, implementing the trained classifiers (e.g., as an extension) into the existing tools (or toolsets) is also worthy of consideration in practice. For example, code review platforms can integrate such classifiers to empower the automated code review process. Such tools may provide more insights and help novices or maintainers troubleshoot architecturally-relevant violation issues.
Suggestion 4: Researchers and practitioners can consider developing plug-ins for code review platforms (e.g., Gerrit) to assist developers in identifying potential architecture violations, or integrating the trained classifiers into existing tools to optimize the automated code review process.
6 Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats by following the guidelines proposed by Wohlin et al. [64].
6.1 Construct validity
Construct validity concerns the connection between operational measures and the studied subjects. The first potential threat to construct validity is the suitability of our selected ML/DL algorithms, feature extraction techniques, and metrics. As mentioned by Peter et al. [65], it is not possible to use an exhaustive approach to test all learning models. We selected five common ML algorithms, one DL algorithm, and three popular pre-trained word embedding models, as they are all widely used to conduct text classification. We admit that using different algorithms and pre-trained word embedding models may generate varying results, despite their popularity. We cannot completely mitigate this threat. In addition, the four metrics (recall, precision, F1-score, and accuracy) are widely used to evaluate the performance of various automatic software engineering techniques (see Section 7.3). Therefore, the metrics pose little threat to the construct validity.
Another potential threat comes from the convenience sampling method of the survey and interview, as individuals with negative views towards the research might be less inclined to participate in our survey and interview. This is a compromise inherent to the convenience sampling method, and it implies that this threat cannot be completely mitigated. Nevertheless, to improve the response rate and collect as many opinions as possible from practitioners, we sent two rounds of reminder emails to the involved practitioners. It is also possible that some survey respondents did not understand the statements well. To reduce this threat to construct validity, we conducted a pilot survey for refining the survey design. Besides, we provided an “I Don’t Know” option in our survey. Moreover, we also interviewed six participants to collect in-depth opinions besides the six defined statements in the survey. To mitigate this threat, we offered an optional open-ended question that allows respondents to freely share their thoughts with us.
6.2 External validity
External validity pertains to the generalizability of our experiment results and findings. We trained the ML and DL-based classifiers with our collected dataset regarding architecture violations. Selecting different datasets and experimental settings might influence the generalizability of (part of) our findings. We acknowledge that the chosen open-source projects are not representative of all projects from the OpenStack and Qt platforms, and our findings may not be generalized to diverse projects, particularly those from different domains such as closed-source projects or projects on other platforms like Apache. Furthermore, the participants of the survey and interviews are contributors to the selected projects, therefore, the findings of RQ3 might not be generalized to other projects. To support further generalizability, we have made our dataset [36] available to enable other researchers and practitioners to replicate, reuse, and extend our study.
6.3 Reliability
Reliability reflects the replicability of a study regarding yielding the same or similar results when other researchers reproduce this study. One potential threat to reliability comes from the experimental settings (e.g., input data and hyperparameters for training classifiers) used in this work. We used the dataset from our previous study [11] that had publicly shared the labeled data and corresponding scripts for collecting data. Besides, we also shared the source code in our replication package [36] to facilitate reproduction and validation. Another potential threat might be from the survey and interview design. To mitigate the threat, we provided the templates of our survey, semi-structured interviews, and customized emails in our replication package [36] to encourage further research in this area.
7 Related Work
In this section, we briefly discuss the current state of research in three associated areas of our study. We first describe related work on code review comments in Section 7.1; we then introduce previous studies on architecture violations in Section 7.2; we finally present the benefits of utilizing ML and DL techniques to analyze artifacts in software development in Section 7.3.
7.1 Code Review Comments
Code review comments contain massive knowledge related to software development, and a variety of studies analyze software defects and evolution through mining review comments and commit records. Han et al. [66] conducted an empirical study by manually checking 25,415 code review comments to investigate code smells identified in modern code review from four OSS projects. Their results show that most reviewers provide constructive suggestions to help developers fix code, and developers are willing to repair the smells through suggested refactoring operations. El Asri et al. [67] empirically investigated the impact of expressed sentiments on the code review duration and its outcome. They found that reviews with negative comments on average took longer time to complete than the reviews with positive/neutral comments. Kashiwa et al. [68] reported Self-Admitted Technical Debt (SATD) during code review by checking 156,372 review records from OpenStack and Qt. They found that 28-48% SATD are introduced during code review, and 20% SATD comments are created because of reviewers’ requests.
Besides, Uchôa et al. [69] investigated the impact of code review on the evolution of design degradation through mining and analyzing a plethora of code reviews from seven OSS projects. They found that there is a wide fluctuation of design degradation during the revisions of certain code reviews. Paixão et al. [70] explored how developers perform refactorings in code review, and they found that refactoring operations are most often used in code reviews that implement new features. Besides, they observed that the refactoring operations were rarely refined or undone along the code review, and such refactorings often contribute to new code smells and bugs.
While the aforementioned studies investigated various aspects of code review regarding development and maintenance (e.g., decisions, design degradation, and SATD), there are no studies that investigated the automated identification of violation symptoms in code review comments; we decided to fill this gap and explore the application of ML/DL techniques to specifically identify violation-related issues (e.g., violation symptoms) from code review comments.
7.2 Architecture Violations
Over the past decades, there have been various investigations on architecture violations. Brunet et al. [71] performed a longitudinal study to explore the evolution of architectural violations in 19 bi-weekly versions of four open source systems. They investigated the life cycle and location of architecture violations over time by comparing the intended and recovered architectures of a system. They found that architectural violations tend to intensify as software evolves and a few design entities are responsible for the majority of violations. More interestingly, some violations seem to be recurring after being eliminated. Mendoza et al. [13] proposed a tool called ArchVID based on model-driven engineering techniques for identifying architecture violations from source code; this tool supports recovering and visualizing the implemented architecture.
Besides, Terra et al. [72] reported their experience in fixing architectural violations. They proposed a recommendation system ArchFix that provides refactoring guidelines for developers and maintainers to repair architectural violations in the module architecture view of object-oriented systems. Their results show that their approach can provide correct recommendations for 75% and 79% of the detected architecture violations on two selected systems, respectively. Maffort et al. [73] proposed an approach to check architecture conformance for detecting architecture violations based on defined heuristics. They claimed that their approach relies on the defined heuristic rules and can rapidly raise architecture violation warnings.
Different from the abovementioned studies focusing on detecting architecture violations in source code, our work explores the possibility of automated identification of the violation symptoms from code review comments.
7.3 Analyzing Software Repositories with Machine Learning and Deep Learning
The application of ML and DL techniques to extract informative knowledge from various artifacts during software development is gaining popularity in the software engineering community. In recent years, a plethora of empirical studies have been conducted to investigate automated techniques for supporting software engineering tasks. Khalajzadeh et al. [53] conducted an empirical study by extracting and manually analyzing app reviews from 12 OSS projects and developed ML and DL models to automatically detect and classify human-centric issues from app reviews and discussions. Their results show that automated techniques can help developers to recognize and appreciate human-centric issues of end-users more easily. Nasab et al. [52] developed 15 ML and DL models to automatically identify security discussions, and then they collected practitioners’ feedback through a validation survey and confirmed the promising applications of the models in practice.
AlOmar et al. [74] proposed an approach to automatically identify and classify self-admitted refactoring in commit messages. They compared their approach with pattern-based and simple random baselines, and the results show that a good performance and a relatively small training dataset is sufficient to classify self-admitted refactoring commits. Hey et al. [75] proposed a novel approach, named NoRBERT, for requirements classification by using the transfer learning capabilities of BERT. They applied their approach to different tasks in the domain of requirements classification, and the results show that their approach improves requirements classification and can be applied to unseen projects with convincing results.
In this work, word embeddings were utilized as word representations and as the input for training ML and DL-based classifiers, as prior studies have demonstrated the effectiveness of word embeddings in capturing rich semantic and syntactic features of words [61]. We employed ML/DL-related techniques to promote the automated identification of violation symptoms. Although much work has been done on automatic text identification, there is currently no automatic way to identify architecture violations in code review comments. Our approach is not restricted to specific textual information, so future studies can replicate and extend our work with other artifacts.
8 Conclusions and Future Work
During code review, reviewers typically spend a substantial amount of effort in comprehending code changes, as significant information (e.g., architecturally relevant information) for inspecting code changes may be dispersed across several files that the reviewers are not acquainted with. automated identification of architecture violations from the discussions between reviewers and developers can save valuable time and effort to locate and check potential issues.
In this work, we attempted to address this issue of automated identification of violation symptoms in code review comments. To this end, we first performed a series of experiments on a dataset of architecture violations from code review comments. Subsequently, to validate the usefulness of our trained ML and DL-based classifiers, we conducted a survey and semi-structured interviews that acquired feedback from the involved developers who discussed architecture violations in code reviews.
Specifically, we developed 15 ML-based and 4 DL-based classifiers to identify violation symptoms from developer discussions in code reviews (i.e., code review comments from four OSS projects in Gerrit). The results show that the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100 and 300-dimensional models. In addition, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperforms the individual classifiers. Moreover, practitioners’ perception of the usefulness of the classifiers’ results validates the promising applications of automated identification of violation symptoms in practice. Especially, such classifiers can enable and inspire practitioners to find architecturally-related violation issues, prioritize, and eventually handle them.
In the future, we plan to conduct a more in-depth investigation of how practitioners deal with architecture violations in code review, what kind of practices and tools they use in dealing with architecture violations, and how using our approach in their daily job can support the identification of architecture violations. We also plan to further enhance the classifiers by identifying fine-grained types of architecture violations based on a larger dataset of more projects from both open-source and industry projects. In particular, a bigger picture is that we plan to delve into a more in-depth investigation and comparison of the performance of trained classifiers as plug-ins or tools with fine-tuned large language models (such as Llama 2) using real-world scenarios. The goal of such plug-ins or tools for code review platforms is to provide warnings of potential architecture violations to practitioners (e.g., developers and reviewers) for them to re-check the related code changes during code review.
Acknowledgements
We would like to thank all the practitioners who participated in our survey and interviews. This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.