### v5

# Towards Automated Identification of Violation Symptoms of Architecture Erosion

Abstract.
Architecture erosion has a detrimental effect on maintenance and evolution, as the implementation deviates from the intended architecture. To prevent this, development teams need to understand early enough the symptoms of erosion, and particularly violations of the intended architecture. One feasible way is through the automated identification of architecture violations from textual artifacts, and particularly code reviews. In this paper, we developed 15 machine learning-based and 4 deep learning-based classifiers with three pre-trained word embeddings to identify violation symptoms of architecture erosion from developer discussions in code reviews. Specifically, we looked at code review comments from four large open-source projects from the OpenStack (Nova and Neutron) and Qt (Qt Base and Qt Creator) communities. We then conducted a survey and semi-structured interviews to acquire feedback from the involved participants who discussed architecture violations in code reviews, to validate the usefulness of our trained classifiers. Moreover, we conducted additional comparative experiments by evaluating Large Language Model (LLM)-based classifiers, including GPT-4o, Qwen-2.5, and DeepSeek-R1. The results show that, for ML/DL-based classifiers, the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Moreover, 200-dimensional pre-trained word embedding models outperform classifiers that use 100- and 300-dimensional models. For ML/DL-based classifiers, ensemble classifiers based on the majority voting strategy can enhance the classifier and outperform the individual classifiers. The findings derived from the online survey and interviews conducted with the involved developers reveal that the violation symptoms identified by our approaches have practical value and can provide early warnings for impending architecture erosion. Furthermore, LLM-based classifiers consistently outperform traditional ML/DL models, with GPT-4o yielding the highest F1-score of 0.851, though ensemble strategies offered no further performance gains for ensemble LLM-based classifiers in our case. We investigate the automated identification of violation symptoms from code reviews using both traditional ML/DL and state-of-the-art LLM techniques. Our contributions include the automated approach for identifying violation symptoms in code reviews, a systematic comparison of ML/DL and LLM approaches, and practitioner‐centered insights into their practical usefulness, ultimately contributing to better architectural conformance and sustainability in software systems.
Architecture Violation, Architecture Erosion Symptom, Natural Language Processing, Machine Learning, Deep Learning, Large Language Models
We would like to thank all the practitioners who participated in our survey and interviews. This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62402348 and 62172311.
journal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software post-development issues
1. Introduction
Software architecture is critical in developing large and complex systems, as it reflects the system structure and behavior and acts as a bridge between business goals and the implemented system (Bass2021sap, ). However, the implemented architectures often diverge from their intended architectures over time, which is the phenomenon of architecture erosion (Li2022SMS, ; PerryWolf1992AEr, ). This phenomenon undermines maintainability, performance, and overall sustainability (Li2022SMS, ; Le2016rad, ; Venters2018ss, ) and often described by various terms in the literature (Li2022SMS, ) and practice (Li2021uae, ), such as architectural decay, degradation, and deterioration. Various symptoms or signs can be observed when architecture erosion occurs during a software system’s life cycle (Li2021uae, ; Li2022SMS, ; Herold2015dvc, ; Fontana2016aer, ; Macia2012anomalies, ). Our recent systematic mapping study (Li2022SMS, ) reported four categories of such architecture erosion symptoms: structural symptoms (e.g., cyclic dependencies), violation symptoms (e.g., layering violation), quality symptoms (e.g., high defect rate), and evolution symptoms (e.g., extensive ripple effects of changes).
In this study, we specifically focus on violation symptoms, as those are the most immediate symptoms of architecture erosion and require the most attention from practitioners (Li2022sae, ; Li2023wvs, ). This symptom type stems from the definition of architecture erosion (Li2022SMS, ), as a violation causes the implemented architecture to diverge from the intended one. Architecturally-relevant violations in software systems include violations of design principles, architecture patterns, decisions, and requirements (Li2022SMS, ). The accumulation of violation symptoms can render the architecture completely untenable (DeSilva2012AErsurvey, ). For brevity, we refer to violation symptoms of architecture erosion as violation symptoms throughout the remainder of this paper.
While a few temporary violation symptoms might be innocuous and might not result in software errors, the accumulation of architecture violations result in architecture erosion (PerryWolf1992AEr, ; DeSilva2012AErsurvey, ; Mendoza2021avd, ); consequently, erosion negatively affects quality attributes, including maintainability and performance (Li2021uae, ; Mendoza2021avd, ). Thus, it is essential to identify and monitor violation symptoms to facilitate the maintenance of the system architecture and further reveal architecture inconsistencies. This is the first step towards, eventually, repairing or mitigating architecture erosion (DeSilva2012AErsurvey, ; Li2022SMS, ).
Violation symptoms can be identified static analysis of source code, but that has certain limitations. As mentioned in previous studies (Sharma2018sss, ; Li2022SMS, ; Li2023wvs, ), a number of factors hinder the effective utilization of static code analysis tools in precisely identifying violation symptoms through source code. For example, such tools only detect limited violation types and narrow programming language support, and cannot capture semantics of violation symptoms by relying solely on code metrics. Alternatively, violations symptoms can be manually identified by analyzing textual artifacts that contain information related to system architecture and design, such as code comments, code reviews, and issue trackers (Li2022sae, ). Recent studies (Li2022sae, ; Li2023wvs, ) have proposed taxonomies on violation symptoms, that can be used as a checklist to detect such symptoms during code review. However, it could be tedious, time-consuming, and potentially error-prone to manually identify potential architecture violations (Li2023wvs, ). Automated techniques could address this problem, but to the best of our knowledge, no prior study has investigated the automated identification of violation symptoms in textual artifacts, such as code review comments.
To address this gap, our goal is to assess the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. The automated identification of violation symptoms could act in a complementary way to modern code review. Modern code review usually addresses code changes in general through knowledge sharing (badampudi2023modern, ), while automated identification of violation symptoms can offer an extra layer of scrutiny. For example, such an automated approach can issue warnings for potential or ignored violation symptoms, preventing the integration of code changes with architectural violations into the code base, and ensuring a more robust and reliable development process. Moreover, automatically identifying violation symptoms from code review comments could offer valuable insights at the architecture level and facilitate the corresponding knowledge sharing between developers and reviewers. The ultimate purpose of our approach is to complement the modern code review process, through warning of and guarding against architectural violation issues, eventually maintaining architectural integrity.
In this work, we first developed 15 classifiers based on Machine Learning (ML) and 4 classifiers based on Deep Learning (DL), to identify violation symptoms from developer discussions in code reviews of four large open-source software projects from the OpenStack and Qt communities. Our results show that the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803. We also found out that utilizing fastText pre-trained word embedding can achieve relatively good results and it helps to improve the classifiers’ performance. Moreover, 200-dimensional pre-trained word embedding models outperform classifiers that use 100- and 300-dimensional models. Finally, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperform the individual classifiers. Then, we conducted an online survey and semi-structured interviews to validate the effectiveness of automated identification of violation symptoms in practice. We received responses from 24 survey participants and 6 interviewees, all of whom were involved in the discussion of the collected review comments concerning architecture violations. Their feedback confirmed the practical value of automated identification of violation symptoms, as such classifiers can support practitioners in finding, prioritizing, and handling architecturally-related issues.
Furthermore, the advent of Large Language Models (LLMs) fundamentally reshaped Software Engineering (SE) paradigms (hou2024large, ; li2025ChatGPT, ), existing studies explored the capacities of LLMs on popular SE tasks (e.g., code generation), but no studies investigate the fine-grained task, identifying violation symptoms in code review comments. Therefore, we further developed LLM-based classifiers using the state-of-the-art LLMs (i.e., GPT-4o, Qwen-2.5, and DeepSeek-R1) and compared their performance with the traditional ML/DL techniques on the same classification task. The results show that LLM-based classifiers typically outperform the traditional ML/DL-based classifiers on all metrics, and GPT-4o-based classifier with 2-shot setting comparatively achieves better performance (F1-score 0.851) than Qwen-2.5 and DeepSeek-R1 on identifying violation symptoms. Notably, ensembling LLM-based classifiers does not improve the performance.
Unlike previous studies that relied on source code analysis to identify architectural issues (see Section 2.1), our work takes a complementary approach by identifying violation symptoms from textual artifacts, which better aligns with practitioners’ perspectives. We also explore ensemble learning techniques to detect architecture erosion symptoms, showing advantages over single ML/DL algorithms. The main contributions of this work are the following:
  • We conducted exploratory experiments to demonstrate the feasibility of ML/DL, LLM-based classifiers for automatically identifying violation symptoms in code review comments, and we compared their performance under different settings, respectively.
  • We investigated the impact of pre-trained word embeddings and shot settings on classification performance. For ML/DL-based classifiers, the SVM-based classifier with a 200-dimensional pre-trained word2vec performs best. Besides, LLM-based classifiers indeed outperform ML/DL-based classifiers, and the GPT-4o-based classifier with a 2-shot setting achieves the best performance.
  • We gained insights from the survey and interview participants regarding the usefulness of trained classifiers in identifying architecture violations during code review. We provide the implications and suggestions based on our empirical study.
The remainder of the paper is organized as follows. Section 2 introduces background information. Section 3 describes the research questions, study design, and experimental setup. The answers to the research questions are presented in Section 4. We discuss the implications of the study results in Section 5 and the threats to validity in Section 6. Section 7 reviews related work, and Section 8 concludes this work with future directions.
2. Background
In this section, we introduce a) the background of architecture erosion and its symptoms; b) architecture conformance checking, the most commonly used approach to detect architecture violations; and c) the modern code review process in Gerrit [^1] https://www.gerritcodereview.com/.
2.1. Architecture Erosion and Related Symptoms
In the past decades, a wide variety of studies are concerned with the architecture erosion phenomenon, which has been extensively discussed and described with various terms (Li2021uae, ; Li2022SMS, ), such as architecture decay (Hassaine2012ADvISE, ; Le2018emad, ), degradation (Herold2020asd, ), degeneration (Hochstein2005cad, ).
Architecture erosion manifests in various symptoms during development and maintenance. A symptom is a partial sign or indicator of the emergence of architecture erosion (Le2016rad, ; Li2022SMS, ; Li2022sae, ). According to a recent mapping study (Li2022SMS, ), erosion symptoms can be classified into four categories (i.e., structural symptoms, violation symptoms, quality symptoms, and evolution symptoms).
Previous studies investigated different symptoms of architecture erosion. Mair et al. (Mair2013tesa, ) proposed a formalization method regarding the process of repairing eroded architecture by finding violation symptoms and recommending optimal repair sequences. Le et al. (Le2016rad, ; Le2018emad, ) regarded architectural smells as structural symptoms and they provided metrics to detect instances of architecture erosion by analyzing the detected smells. Martin (Martin2000dpdp, ) deemed that the evolution symptoms of architecture erosion include rigidity (a tendency to resist changes), fragility (a tendency to break frequently when modifications occur), immobility (inability to be reused), and viscosity (reduced effectiveness and efficiency due to design or environment issues).
Table 1. Examples of violation symptoms collected from Nova, Neutron, Qt Base, and Qt Creator
Project Example of violation symptoms of architecture erosion
Nova But here don’t we have to make a upcall from compute to api db, which will violate api/cell isolation rules. Is there any workaround in this case?
We could get some race conditions when starting the scheduler where it would not know the allocation ratios and would have to call the computes, which is a layer isolation violation to me.”
Neutron Having said all of that: I get that I’m violating an abstraction layer in LinkLocalAddressPair and that this is surprising (and therefore bad).
As you pointed out, since this patch is about the router availability zone it seems like a layer violation doing it in the mech driver.”
Qt Base I think it’s a layer isolation violation to just make the call here. It should be fixed at the Compute API level rather IMHO.
Feels like this DB work violates the level of abstraction we are expecting here.
Qt Creator The unit tests are designed to not depend on any qt creator library to make the dependency breaking easier. Sometimes it works but it is not designed that way.
This looks like we break the design. Why do we can not hold a pointer to the plugin?
As mentioned in the Introduction section, we focus on the most direct erosion symptoms of architecture erosion (i.e., violation symptoms). Table 1 presents several examples of violation symptoms of architecture erosion from the used dataset (see Section 3). To facilitate the examination of additional examples, we have made our dataset publicly available online (onlinepackage, ); a detailed comparison of our work and related work is given in Section 7.
2.2. Architecture Conformance Checking
Architecture Conformance Checking (ACC) techniques are the most commonly-used approaches to detect architecture violations (Li2022SMS, ). They either perform static or dynamic analysis, by comparing the structure of the intended architecture (provided by the architects) with the extracted architecture information from source code. For example, Miranda et al. (Miranda2016acc, ) presented an ACC approach based on static code analysis techniques with a supporting tool ArchRuby. Their tool provides means to control the erosion process by reporting and visualizing architectural violations in two high-level architectural models, namely reflexion models and Dependency Structure Matrices (DSMs).
In addition, rule-based conformance checking approaches are also employed to identify architecture violations. For example, previous studies identified architecture violations by checking explicitly defined architectural rules (Mendoza2021avd, ). Moreover, it is viable to check architecture conformance and identify architecture violations by defining and describing the systems through Architecture Description Languages (ADLs), or Domain-Specific Languages (DSLs).
Despite their potential, the previously mentioned approaches for identifying architecture violations have clear limitations. For instance, much effort is required to address the challenges of understanding the architecture design (e.g., concepts and relations), defining architectural rules (or description languages) in advance, and establishing a mapping between architectural elements and source code (Li2022SMS, ; Li2023wvs, ). Additionally, other limitations, such as a lack of generalizability and architectural views visualization, insufficient tooling support, have impeded the wide adoption of these approaches in practice (Li2022SMS, ).
During software development, textual artifacts (e.g., code review comments) contain rich and valuable information regarding architectural design and changes apart from source code. Mining architectural violations from textual artifacts can significantly facilitate code comprehension for software engineers (Li2023wvs, ). In contrast to the previously mentioned approaches, our study focuses specifically on mining violation symptoms from textual artifacts.
2.3. Code Review in Gerrit
Code review is a crucial software development activity that involves the systematic examination of assigned code to identify defects and improve software quality (Bacchelli2013eoc, ). A methodical code review process not only enhances the quality of software systems, but also facilitates the sharing of development knowledge and prevents the release of unstable and defective products. Over time, code review practices have become increasingly important, have been widely adopted in modern software development, and have received extensive tool support. In fact, tool-based code review has become the norm in both industry and open-source communities, with a variety of available code review tools, such as Gerrit, GitHub’s Pull Requests, Me’s Phabricator [^2] https://www.phacility.com/, and VMware’s Review-Board [^3] https://www.reviewboard.org/.
Figure 1. Code review process in Gerrit
The two communities (i.e., OpenStack and Qt), from which the four open-source software (OSS) projects were selected in this work, use Gerrit as the code review tool. Figure 1 illustrates the modern code review process in Gerrit, which is a well-known code review tool and widely used in modern code review (Davila2021SLRmcr, ). Gerrit allows for efficient collaboration on code changes before they are merged into the main code base. Once a developer commits new code changes (e.g., software patches) and their descriptions to Gerrit, the system creates a page to record all changes and relevant descriptions (e.g., commit messages). Then, the system conducts a sanity check to verify whether the patch is compliant and to ensure that the code does not contain obvious compilation errors. After the submitted patch passes the sanity check, code reviewers manually examine the patch and provide comments to correct any potential errors, followed by giving a voting score. The comments are subsequently used by the developers to improve the patch and submit it for another review. Finally, the submitted patch is merged into the code repository only after passing integration tests (potentially over multiple iterations), which confirms that there are no issues or conflicts. Alternatively, the patch can be abandoned if the voting is negative.
The above process is used, among others, in revisions that enable the analysis of potentially significant architectural changes. Code review comments provide developers a finer granularity for investigating architectural changes and violations in their daily development routines (Li2021mpl, ). Meanwhile, developers provide reasoning and rationale for the changes they make in code review comments, which enables code review data to be a valuable source of knowledge for explaining changes (Paixao2021crac, ), offering insights into developers’ concerns, suggestions, and potential violations.
3. Study Design
This section describes the research questions that motivated this research, followed by a detailed explanation of the data collection and the ML/DL models to automatically identify architecture violations from textual artifacts. Then, we provide details of the validation survey and interview conducted to assess the effectiveness of automated identification of violation symptoms in practice. Furthermore, we introduce and compare the state-of-the-art LLMs on the same task with traditional ML/DL approaches.
Figure 2 provides an overview of the study design. As illustrated in this figure (see the top box, Dataset Creation), we collected data and established a dataset on architecture violations in our previous work (Li2023wvs, ), which contains code review comments related to architecture violations During Phase 1, we explored the possibility of using this dataset to train ML/DL-based classifiers to automatically identify violation symptoms in code review comments. During Phase 2, we sent the survey questionnaire to 200 involved participants and received 24 responses that were used to validate the usefulness of automated identification of violation symptoms in practice. During Phase 3, we constructed LLM-based classifiers through prompts using state-of-the-art LLMs and compared their performance against the ML/DL methods from Phase 1.
Figure 2. An overview of the research process
3.1. Research Questions
In this work, we aim to explore the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. Specifically, we formulated the goal of this study by following the Goal-Question-Metric method (Basili1994gmq, ): analyze code review comments for the purpose of exploring the feasibility of automatically identifying violation symptoms of architecture erosion with respect to the performance on architecture erosion identification from the point of view of developers in the context of open source software development. To achieve our goal, we formulate four Research Questions (RQs):
RQ1: How accurately can classifiers identify violation symptoms from code review comments? RQ1.1: Which classifier performs best in identifying the violation symptoms? RQ1.2: Which word embedding models can help to improve the performance of classifiers? RQ1.3: To what extent, does the dimensionality of word embedding models impact the performance of classifiers?
Rationale: This RQ focuses on training feasible classifiers to automatically identify violation symptoms from textual artifacts. The percentage of violation symptoms (belonging to architecture-level issues) is relatively lower than code-level issues. We thus expect that the description of violation symptoms is scattered in textual artifacts during development. RQ1.1 focuses on finding out the classifier with the best performance to distinguish violation symptoms from natural language by comparing several popular ML/DL algorithms. RQ1.2 aims at employing several well-known word embedding models to generate word vectors as features of labeled data, in order to choose the word embedding models that can help to improve the accuracy of classifiers. The word embedding models are a common ground for both ML/DL-based classifiers in our study. Finally, with RQ1.3, we seek to compare the differences in dimension size among different word embedding models, as it is one of the most important hyperparameters affecting the model performance(hutter2019aml, ). Specifically, RQ1.3 endeavors to shed light on the extent to which the selection of dimensionality impacts the performance of the classifiers in automatically identifying violation symptoms.
RQ2: To what extent, a voting strategy can help to improve the performance of identifying violation symptoms from code review comments?
Rationale: After training individual classifiers based on ML/DL algorithms, we want to know whether, and to what extent, an ensemble classifier composed of trained classifiers can help to improve the performance of identifying violation symptoms from code review comments. To answer this RQ, we compared the performance of individual classifiers with that of a combined classifier. More specifically, we employed a voting strategy to integrate the prediction values of each classifier and generated results after voting among individual classifiers. Answering this RQ can help to explore the possibility of further improving the performance of automated identification of violation symptoms.
RQ3: Do practitioners find automated identification of violation symptoms useful in practice?
Rationale: If the results generated by the classifiers based on ML/DL algorithms are acceptable regarding the effectiveness of identified violation symptoms, automatically identifying violation symptoms from textual artifacts would be helpful in practice. Such automated techniques could facilitate tasks such as locating architecture violations and inspiring further investigation and remediation of similar architectural issues. To answer this RQ, we conducted a survey and semi-structured interviews to investigate how practitioners (both developers and reviewers) assess the effectiveness and applicability of the trained classifiers.
RQ4: Do state‑of‑the‑art LLM-based classifiers truly outperform traditional ML/DL classifiers on the same classification tasks, and if so, to what extent do they offer performance gains?
Rationale: Considering the remarkable performance of LLMs on various SE tasks (hou2024large, ; li2025ChatGPT, ), we aim to further validate our conjecture whether LLM-based classifiers truly outperform traditional ML/DL-based classifiers regarding the automated identification of violation symptoms in code review comments. To answer RQ4, we developed several LLM-based classifiers using different prompt strategies to perform the same classification tasks in RQ1 and RQ2, enabling a fair comparison of performance gains.
3.2. Data Collection
We relied on the dataset that we mined and manually identified in our previous study (Li2023wvs, ). As shown in Figure 2 (see Dataset Creation), we followed five steps to establish a specific dataset on architecture violations mined from code review comments. More specifically, we mined code review comments of four OSS projects in the OpenStack (i.e., Nova and Neutron) and Qt (i.e., Qt Base and Qt Creator) communities between 2014 and 2020. We mined the data through the REST API [^4] https://gerrit-review.googlesource.com/Documentation/rest-api.html supported by Gerrit. We first conducted experiment-based keyword refinement to improve our keyword search, as our previous experience indicated that a keyword-based approach could only capture very small proportions of code review comments containing violation symptoms (Li2023wvs, ). The dataset collection process comprised three main steps: keywords searching, semantically-similar words querying (identifying potentially semantic-related data), and co-occurrence matrix-based querying (capturing potentially co-occurred and associated data). The whole manual labeling and analysis process took the researchers around one and a half months. Subsequently, we proceeded with manual identification, and pilot and formal data labeling of review comments related to architecture violations. In total, we manually identified and labeled 606 code review comments pertaining to architecture violations. Code review comments were created for patches, and comments were made by code reviewers and developers. In terms of the data labeling process, four researchers first conducted a pilot data labeling by randomly selecting 50 review comments, to reach a consensus (the inter-rater agreement between the researchers measured in the Cohen’s Kappa coefficient value (Cohen1960cans, ) was 0.857) and ensure that we have the same understanding of violation symptoms. Any disagreements were discussed between the four researchers to reach a consensus. After that, the four researchers started the formal data labeling by dividing the retrieved 21,583 code review comments into four parts (each researcher manually labeled around 5,400 comments). After the formal data labeling, the first author checked the data labeling results from the other three researchers. Meanwhile, to mitigate potential bias, the four researchers discussed all the conflicts in the labeling results until we reached an agreement. To summarize, all data labeling results were checked by at least two researchers and conflicts were discussed among four researchers. Subsequently, all the collected code review comments regarding architecture violations (i.e., 606 code review comments) from the four projects were combined into an integrated dataset used in this work. The dataset is available in our replication package (onlinepackage, ).
3.3. Phase 1 - Automated Classification of Violation Symptoms based on ML/DL Techniques
In this section, we describe the experimental setup that we followed to evaluate the performance of the trained classifiers (see Figure 3). The steps in this setup correspond to the steps of Phase 1 in Figure 2. First, we pre-processed the collected code review comments to generate structured word sets, and then utilized word embedding techniques to encode the words and generate vectors for presenting the words in review comments. Subsequently, the generated vectors acted as the input to train the classifiers based on ML/DL models, in order to learn to classify review comments as violations and non-violations. The experimental environment is a computer equipped with Intel(R) Core(TM) i7-10510U CPU and 16GB RAM, running Windows 11 (64-bit). Strictly speaking, DL is a subset of ML. However, to clearly differentiate between traditional ML and DL approaches, we provide a detailed explanation of these steps for ML/DL models separately in the following subsections. More details of the experiments (such as hyperparameters and experimental environment) can be found in the replication package (onlinepackage, ).
Figure 3. The framework of the experimental setup for classifying violation symptoms of architecture erosion from code review comments
3.3.1. Machine Learning Models
Before training ML models, we need to pre-process our collected data, which involves transforming raw data into a structural format for model training. The pre-processing includes five steps listed as follows:
Step 1. Tokenization: The process of tokenization is to break a stream of text into words, punctuation, and other meaningful elements called tokens. In this work, we tokenized the review comments by splitting the text into its constituent set of words.
Step 2. Noise Removal: Developers usually use plain text when they discuss violation symptoms during code review, which contain various noise data such as punctuation, numbers, and special characters (e.g., $\backslash$”, “*”). Noise data usually does not contain valuable semantic information, and we therefore removed it.
Step 3. Stop words Removal: Stop words are frequently observed in text, for the purpose of distinguishing between texts, such as “the”, “are”, and “is”, but they are typically devoid of valuable semantic content. Thus, we removed stop words in our dataset by using the Natural Language Toolkit (NLTK) (Bird2010nlp, ).
Step 4. Capitalization Conversion: In order to ensure the consistency of word form and to avoid redundant counts of words, all textual data was converted to lowercase.
Step 5. Stemming: Stemming can be used to convert the words to their base forms. For instance, “architecture” and “architectural” have the same stem (i.e., token) “architectu”. We conducted a stemming process (using Snowball Stemmer from the NLTK toolkit (Bird2010nlp, )) to obtain the stem of each token, since it is a more comprehensive and aggressive version of the Porter Stemmer (SnowballStemmer, ).
Word Embedding. Feature extraction techniques are widely utilized to transform unstructured text sequences into structured feature spaces (e.g., matrices, vectors, or encodings). In this work, we employ word embedding techniques to extract textual features for generating interpretable input for machine learning algorithms. Word embedding is a prominent method employed in natural language processing to represent words through dense vectors. In this work, we utilized the three prevalent and popular techniques (i.e., word2vec, fastText, and GloVe) to generate word embeddings. These techniques have gained widespread use in machine learning and deep learning-based tasks, owing to their efficacy in capturing the semantic and syntactic relationships among words.
  • word2vec: word2vec was developed in 2013 (Mikolov2013word2vec, ) and it takes words as input and provides a word embedding matrix that contains high-dimensional unique vectors to represent each distinct word in given training corpora. word2vec can generate word embeddings through skip-gram and Continuous Bag-of-Words (CBOW) models. In this study, we adopted the pre-trained word2vec model proposed by Efstathiou et al. (Efstathiou2018weso, ). This model was trained based on the corpus comprising over 15 GB of textual data collected from Stack Overflow posts, which contains a plethora of textual information in the SE domain.
  • fastText: fastText is an open-source and lightweight library proposed by Meta AI Research lab in 2017 (Bojanowski2017fastText, ) and it is capable of learning text representations and training text classifiers. Based on the fastText model, Meta published pre-trained word vectors with 300 dimensions for 157 languages constructed on Common Crawl and Wikipedia [^5] https://fasttext.cc/docs/en/crawl-vectors.html. In this study, we employed the pre-trained fastText model, which can be transformed into 100 and 200-dimensional models, to control variables and avoid the impact of dimensionality on classifiers’ performance.
  • GloVe: Global Vectors for Word Representation (GloVe) is another common word embedding technique (Pennington2014GloVe, ). A GloVe model is inspired by aggregated word co-occurrence statistics that may encode global information for words. Compared with word2vec, GloVe avoids the weakness regarding using local word co-occurrence information. In this study, we adopted the GloVe pre-trained word vectors on Twitter data (two billion tweets) with 200 dimensions.
Model Training. Given the binary classification task in this study, we selected five common ML algorithms that are widely used in classification tasks (Han2022dm, ), including Support Vector Machine (SVM), Logistic Regression (LR), Decision Tree (DT), Bernoulli Naive Bayes (NB), and k-Nearest Neighbor (kNN). These algorithms are also commonly applied in areas of SE studies (Yang2022pmse, ).
For the ML classifiers, we performed a grid search to tune the hyperparameters of each model and picked the best-performing ML models with specific hyperparameters on the validation set. Grid search, an exhaustive search technique, explores every possible combination within a predefined search space to identify the optimal configuration. Specifically, it is widely employed in the realms of machine learning and optimization and it involves a systematic evaluation that examines all of the combinations of a set of candidate settings to find the best hyperparameter combination (hutter2019aml, ).
3.3.2. Deep Learning Models
Similarly to the ML models, for the DL models we pre-processed the collected data before training DL classifiers, using the same pre-processing steps described in Section 3.3.1. For the pre-processing of DL classifiers, we needed to ensure the equal size of each paragraph length; therefore, we utilized the zero-padding strategy before dimension transformation to equalize all input text (i.e., the maximum length is set to 2000 words).
Model Training. In recent years, DL-based models are increasingly used in various text classification tasks and Convolutional Neural Network (CNN) has become one of the most popular model architectures for text classification (Minaee2021dltc, ). We selected the TextCNN model to train the DL-based classifiers, since TextCNN, as an adapted CNN-based model, is reported to improve the state of the art on text classification (Minaee2021dltc, ). TextCNN is a CNN-based architecture designed for processing textual data, and was introduced by Yoon Kim in 2014 (Kim2014TextCNN, ). It can be used for a variety of Natural Language Processing (NLP) tasks, such as text classification and sentiment analysis. Compared with traditional ML models regarding text processing, the key advantage of TextCNN is its ability to automatically learn features from raw textual data, without relying on manual feature engineering.
We built the TextCNN for identifying violations and non-violations in code review comments through several processing layers, including an input layer, a convolutional layer, a pooling layer, a fully connected layer, and an output layer. Specifically, the TextCNN model applies one layer of convolutions over n-dimensional word embeddings to capture n-gram features at different scales. The convolutional layer learns to detect local patterns in the input text. The resulting feature maps are then fed through one max-pooling layer, which extracts the most salient features from each map and concatenates them to produce a fixed-length vector representation of the input text. This vector is then passed through one or more fully connected layers to produce the final output, i.e., the predicted labels for the input review comments.
TextCNN is a convolutional neural network model that is commonly used for text classification tasks. This model takes a sentence as input and applies multiple convolutional filters to extract important features from the text. A TextCNN classifier based on the vocabulary set of a dataset uses a one-hot encoding scheme to represent words and generate the word embedding, while a TextCNN classifier can also utilize pre-trained word embedding models to capture semantic information of words. Additionally, the TextCNN model requires a proper setting to construct neural network models, such as hyperparameters and word embedding parameters. In this work, we set the embedding dimension to 200, the learning rate to 0.001, and the dropout to 0.25 to handle the overfitting, while the filter region size is (3, 4, 5). Each DL model’s batch size is set to 16, and the network is trained for 100 epochs with the early stopping strategy with the purpose of avoiding overfitting. We configured the patience parameter to 8, indicating that the training process would terminate after 8 consecutive epochs with no observed improvement in model performance. These approaches and settings are common practices in work related to ML/DL and text classification, aimed at preventing overfitting and optimizing the model’s convergence (Watson2022SLRDL, ).
3.3.3. Performance Evaluation
For the performance evaluation of ML/DL classifiers, we adopted four commonly used metrics, namely, Precision, Recall, F1-score (i.e., the harmonic mean of Precision and Recall), and Accuracy.
Precision presents the proportion of correctly classified violation symptom comments to all comments classified as violation symptoms. Recall represents the proportion of all violation symptom comments that are correctly identified. F1-score is the harmonic mean between precision and recall, assessing if an increase in one compensates for a decrease in the other. Accuracy is the percentage of correct classifications by a trained learning model. The above metrics are defined and calculated by the following equations.
(1) $\rm Precision=\frac{TP}{TP+FP}$
(2) $\rm Recall=\frac{TP}{TP+FN}$
(3) $\rm F_{1}-score=2\times\frac{Precision\times Recall}{Presion+Recall}$
(4) $\rm Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
Here, we need to use four statistics to calculate the metrics, namely, False Positives (FP) represent the number of non-violation comments that are classified as violation comments; False Negatives (FN) represent the number of violation comments that are classified as non-violation comments; True Positives (TP) represent the number of violation comment that are classified as violation comments; True Negatives (TN) represent the number of non-violation comments that are classified as non-violation comments.
To avoid the class imbalance problem (he2009imbalance, ; Buda2018SMSCNN, ), a data distribution problem that significantly impacts classifiers’ performance, we constructed a balanced dataset containing equal numbers of review comments labeled as “violation” and “non-violation”, respectively. Note that, we randomly selected 606 “non-violation” comments to balance the dataset, and this balanced distribution mitigates the poor generalization ability of classifiers, and avoids leading to highly misleading assessment (Buda2018SMSCNN, ). With respect to the model training and validation strategy, we employed 10-fold cross-validation when we conducted a performance evaluation of the classifiers. We divided the collected dataset into three parts, i.e., we randomly set aside 80% of the data for the training set (60%) and validation set (20%), and 20% for the test set. Besides, we set a fixed seed for reproducibility during experiments.
3.3.4. Ensemble Classifier
To identify architecture violations from textual artifacts (i.e., code review comments), we trained several ML/DL-based classifiers to predict the labels of new review comments. However, the trained classifiers have varying performances in predicting the labels of target artifacts. In RQ2, we asked whether composing the individual classifiers into an ensemble classifier can help improve the overall performance. This composition can be done by utilizing the ensemble learning technique, i,e., majority voting strategy (Dietterich2000emml, ). Specifically, a combined classifier of several individual classifiers is generated with the purpose of building an improved classifier, which predicts an output (class) based on the highest probability of chosen class as the output (Dietterich2000emml, ). This aggregating criterion is like an election process and can produce combined voting results from each classifier’s output.
The voting criteria include Hard Voting (voting is calculated on the predicted output class) and Soft Voting (voting is calculated on the predicted probability of the output class). Considering the computational and memory requirements, we employed the Hard Voting strategy (a cost-efficient majority voting strategy) in this work. In other words, the final predicted results depend on the majority output (i.e., labels) of the individual classifiers.
3.4. Phase 2 - Validation: Survey and Interview
To answer RQ3, we designed and executed a survey and semi-structured interviews by following the guidelines proposed by Kitchenham and Pfleeger’s personal opinion surveys (gaese2008CH3, ); in terms of the survey, we used an anonymous, web-based questionnaire to increase the response rate.
(1) Survey. Surveys can be conducted in different ways, such as online questionnaires, phone surveys (Lethbridge2005sse, ). Our survey is an online questionnaire of the cross-sectional type, which allows us to collect data and compare differences between participants at a single point in time. We adopted self-administered questionnaires to collect responses, since (a) respondents can answer the survey at their convenience (reducing their time investment and increasing the response rate); (b) questionnaires can reduce research bias since the researcher does not have any contact with the respondents when they fill out the survey; and (c) respondents enjoy better privacy than interviews.
We invited all the developers involved in our collected dataset through customized emails (see the replication package (onlinepackage, )) attaching the links to their code review comments that contain violation symptoms and are identified by our approach (i.e., the best-performing classifier), as well as the survey form (also see the replication package (onlinepackage, )). Our survey was constructed based on four parts: Problem Statement, Our approach, Survey Goal, and Survey Questions. Problem Statement introduces the research background regarding what is architecture erosion and its violation symptoms, as well as why it is important to investigate violation symptoms. Our approach describes the approach that we developed to automatically identify violation symptoms with high performance from a large number of code reviews as well as two examples of violation symptoms of architecture erosion detected by our approach from code review comments. In Survey Goal, we explain the purpose of this survey. In Survey Questions, we ask five general questions and present six statements concerning the usefulness of our approach. Note that, before the formal survey, we did a pilot survey (receiving 6 responses) with the purpose of revising and improving our survey questions and examples.
The five general survey questions include demographic information (e.g., their countries and work experience). The six statements are assessed using a five-point Likert scale (Strong Agree, Agree, Neutral, Disagree, and Strongly Disagree) with one option (“I Don’t Know”). This scale enables participants to rate the extent that they agree or disagree with the statements. At the end of the survey, we ask an open-ended question to let the respondents freely provide comments and suggestions about our survey. To formulate the statements, we took inspiration from the survey questions of two survey studies (Nasab2021ais, ; Khalajzadeh2022sda, ) that evaluate the usefulness of automatic classifiers based on ML/DL techniques in the SE community. We chose to use declarative statements in our validation survey since declarative statements are deemed suitable for exploratory studies to convey simple and clear judgments; in contrast, descriptive statements are more suitable in explanatory studies to provide detailed judgments in various aspects. The statements are as follows:
  • Statement 1. “Violation symptoms in code review comments identified by the approach, represent potential architectural violations”.
  • Statement 2. “Violation symptoms in code review comments identified by the approach can be used to improve the code quality”.
  • Statement 3. “I (as a practitioner) can find potentially constructive and useful architectural information from violation symptoms in code review comments identified by the approach”.
  • Statement 4. “Violation symptoms in code review comments identified by the approach can help us identify violation-related issues faster than if we do this manually”.
  • Statement 5. “Violation symptoms in code review comments identified by the approach might help us locate, identify, and prioritize potential architectural violations in our systems”.
  • Statement 6. “Violation symptoms in code review comments identified by the approach can provide us with input to find other violations of similar nature or otherwise”.
(2) Interview. As a supplement to the survey, we also invited the same practitioners to participate in semi-structured interviews for more in-depth discussions. Specifically, the interview focused on the developers’ opinions on classifier effectiveness, as well as the role of such automated techniques in the code review process during software development. The semi-structured interviews were conducted by the following three open-ended Interview Questions (IQs):
  • IQ1: How do you think the integration of an automated system for identifying violation symptoms of architecture erosion from code review comments would fit within the code review workflow? This system would analyze reviewer comments in real-time to flag potential or overlooked violation symptoms. By doing so, it would prevent the incorporation of code changes that violate system architecture, thereby contributing to a more robust and reliable development process. What are your thoughts on the potential benefits and challenges of such integration? This question aims to assess the practical feasibility and acceptance of integrating automated violation symptoms identification into existing code review workflows.
  • IQ2: Based on your experience, what strategies would you recommend to effectively use automated identification of violation symptoms in the code review process? How can this technology be utilized to maximize its benefits and efficiency in identifying and addressing architectural issues during code review? This question seeks to gather expert insights on optimal deployment strategies and best practices for automated violation symptoms identification.
  • IQ3: In addition to code review scenarios, can you envision other contexts or scenarios where our approach for automated identification of violation symptoms could be beneficial in software development practice? What potential applications or integrations do you see as valuable in enhancing code quality and development workflows? This question explores the broader applicability and extensibility of automated violation symptoms identification beyond code review contexts.
Participants. Our collected data from Gerrit contains the developers’ information (i.e., names and email addresses); this information is not included in the replication package for privacy reasons. For the survey, we invited the potential participants to fill out our survey by sending emails to all developers whose review comments are included in our collected dataset. During the pilot survey, we got responses from 6 developers. The pilot survey helped us to evaluate and refine our survey design. For example, we added more precise examples of violation symptoms from code review comments in our survey in order to enhance practitioners’ understanding, and we refined the content of the customized emails based on the received pilot feedback. Then, we invited the remaining developers involved in the discussions on violation symptoms to fill in the survey. Finally, for the semi-structured interviews, we invited the above-mentioned participants to give feedback on our trained classifiers and the role of such automated techniques during code review. We first conducted two pilot interviews and found that, the interview questions are suitable. Subsequently, we interviewed the remaining participants who accepted our invitation, resulting in a total of six interviews.
Sample and Population. The target population is limited to developers whose code review comments are identified by our approach. We used a Non-Probabilistic Sampling method in this survey, namely, Convenience Sampling. As a type of non-probability sampling method, Convenience Sampling can help conveniently reach a suitable number of developers who are willing to participate in our survey. The reason we chose this sampling method is that we could not intentionally choose a sample due to the limitations of geographical constraints and scheduling conflicts.
Data Analysis. For the survey questions, we applied descriptive statistics to analyze the responses to the demographic and Likert scale questions. For the open-ended interview questions, interviews were recorded and transcribed; subsequently, we used the Open Coding and Constant Comparison method (Adolph2011CC, ) (a fundamental qualitative analysis technique widely used in empirical software engineering research) to analyze and categorize the qualitative responses.
3.5. Phase 3 - Comparison: LLMs vs. ML/DL Techniques
Unlike traditional ML/DL approaches, LLMs eliminate the need for task-specific training from scratch. Therefore, we leveraged pre-trained and open-source LLMs using techniques like prompt engineering to perform the same classification tasks in Phases 1 and 2. First, we conducted preliminary experiments to identify high-performing prompts and adopted the best-performing prompt for subsequent formal evaluations. We employed both zero-shot and few-shot learning approaches, explicitly evaluating the few-shot setting with 0, 2, 4, 6, and 10 examples, and assessed how performance varied across three representative LLMs: GPT-4o, Qwen-2.5, and DeepSeek-R1.
LLM Selection. We selected three representative and high-performance LLMs based on their technical diversity, capabilities, and availability.
  • GPT-4o: A proprietary multimodal LLM developed by OpenAI (GPT-4o, ). Although its architecture is undisclosed, industry speculation suggests that it involves approximately 200 billion parameters. GPT-4o demonstrates state-of-the-art performance in code generation (HumanEval) and mathematical reasoning (MATH), and supports real-time multimodal interactions with an average response latency of 320 ms. Primary applications include enterprise AI, educational systems, and multimedia content generation.
  • Qwen-2.5: An open-source LLM from Alibaba’s DAMO Academy with 72 billion parameters. Qwen2.5-72B (qwen2.5, ) supports a 128k-token context window and excels in multilingual performance, particularly in Chinese, as demonstrated by the C-Eval benchmark. It supports 12 languages and achieves ChatGPT-4-level comprehension.
  • DeepSeek-R1: An open-source 671-billion-parameter model (deepseekai2025, ) using a Mixture of Experts (MoE) design with only 8 billion active parameters per inference. This architecture offers a 60% reduction in training cost and over 2.3× inference throughput compared to dense models like LLaMA 3-70B. Its efficient MoE architecture and strong reasoning capabilities make it particularly suitable for evaluating performance across diverse SE tasks while maintaining computational efficiency.
Zero-Shot and Few-Shot Learning. LLMs inherently support zero-shot and few-shot learning due to their robust generalization and natural language understanding capabilities.
  • Zero-shot learning involves presenting no labeled examples. Instead, the model receives only a conceptual description (e.g., “violation symptoms”) and must perform classification based on this abstract guidance alone.
  • Few-shot learning supplements the conceptual definition with a small set of labeled examples. The model uses these demonstrations to infer classification strategies, blending prior knowledge with task-specific cues.
To investigate performance variability, we tested four different few-shot settings (2, 4, 6, and 10 examples), ensuring class balance by selecting even numbers of samples.
Prompt Engineering. Prompt engineering is critical in guiding LLMs toward accurate and reliable outputs. Properly designed prompts significantly influence performance, particularly in complex classification tasks. While optimal prompt design is often non-trivial due to the large design space, we adopted an empirical evaluation approach. We generated multiple prompt candidates for both zero-shot and few-shot learning and evaluated their performance in preliminary trials, and then selected the best-performing prompt. The prompt defined the LLM’s assumed role, introduced the concept of violation symptoms, provided demonstrations (in few-shot settings), and specified the required output format.
4. Results
4.1. RQ1: Identifying Violation Symptoms
To answer RQ1, we trained five ML classifiers and one DL classifier based on three word embedding techniques (i.e., word2vec, fastText, and GloVe). Moreover, we evaluated the performance of these classifiers and compared the differences between their performance. In general, we use F1-score as the overall evaluation of classifiers’ performance.
Table 2. Performance comparison among ML-based classifiers
Classifier Precision Recall F1-score Accuracy
w2v FT GloVe Average w2v FT GloVe Average w2v FT GloVe Average w2v FT GloVe Average
SVM 0.789 0.728 0.759 0.759 0.828 0.746 0.828 0.801 0.808 0.737 0.792 0.779 0.803 0.734 0.783 0.773
LR 0.832 0.777 0.798 0.802 0.730 0.713 0.713 0.719 0.777 0.744 0.753 0.758 0.791 0.754 0.766 0.770
NB 0.728 0.743 0.717 0.729 0.615 0.615 0.623 0.618 0.667 0.673 0.667 0.669 0.693 0.701 0.689 0.694
DT 0.667 0.657 0.643 0.656 0.689 0.721 0.664 0.691 0.677 0.688 0.653 0.673 0.672 0.672 0.648 0.664
kNN 0.863 0.788 0.797 0.816 0.516 0.549 0.451 0.505 0.646 0.647 0.576 0.623 0.717 0.701 0.668 0.695
Average 0.776 0.739 0.743 0.753 0.676 0.669 0.656 0.667 0.715 0.698 0.688 0.700 0.735 0.712 0.711 0.719
Table 3. Performance comparison among DL-based classifiers
Classifier Precision Recall F1-score Accuracy
TextCNN_voc 0.475 0.541 0.469 0.471
TextCNN_w2v 0.584 0.680 0.596 0.598
TextCNN_FT 0.728 0.549 0.667 0.672
TextCNN_GloVe 0.507 0.885 0.434 0.512
Average 0.574 0.664 0.542 0.563
Table 4. Performance comparison of the classifiers based on the fastText model with different dimension size
Classifier Precision Recall F1-score Accuracy
100-dim 200-dim 300-dim 100-dim 200-dim 300-dim 100-dim 200-dim 300-dim 100-dim 200-dim 300-dim
SVM 0.745 0.728 0.657 0.672 0.746 0.754 0.707 0.737 0.702 0.721 0.734 0.680
LR 0.704 0.777 0.745 0.664 0.713 0.648 0.684 0.744 0.693 0.693 0.754 0.713
NB 0.745 0.743 0.712 0.623 0.615 0.607 0.679 0.673 0.655 0.705 0.701 0.680
DT 0.641 0.657 0.612 0.615 0.721 0.648 0.628 0.688 0.629 0.635 0.672 0.619
kNN 0.831 0.788 0.831 0.525 0.549 0.443 0.643 0.647 0.578 0.709 0.701 0.676
TextCNN 0.577 0.728 0.494 0.648 0.549 0.730 0.585 0.667 0.461 0.586 0.672 0.492
Table 5. Performance comparison of ensemble ML/DL-based classifiers
Classifier Precision Recall F1-score Accuracy
Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B
SVM 0.759 0.789 0.795 4.74% 0.76% 0.801 0.828 0.828 3.37% 0.00% 0.779 0.808 0.811 4.11% 0.37% 0.773 0.803 0.807 4.40% 0.50%
LR 0.802 0.832 0.833 3.87% 0.12% 0.719 0.730 0.738 2.64% 1.10% 0.758 0.777 0.783 3.30% 0.77% 0.770 0.791 0.795 3.25% 0.51%
NB 0.729 0.743 0.738 1.23% -0.67% 0.618 0.615 0.623 0.81% 1.30% 0.669 0.673 0.676 1.05% 0.45% 0.694 0.701 0.701 1.01% 0.00%
DT 0.656 0.657 0.689 5.03% 4.87% 0.691 0.721 0.746 7.96% 3.47% 0.673 0.688 0.717 6.54% 4.22% 0.664 0.672 0.705 6.17% 4.91%
kNN 0.816 0.863 0.836 2.45% -3.13% 0.505 0.516 0.500 -0.99% -3.10% 0.623 0.646 0.626 0.48% -3.10% 0.695 0.717 0.701 0.86% -2.23%
TextCNN 0.574 0.728 0.537 -6.45% -26.24% 0.664 0.549 0.713 7.38% 29.87% 0.542 0.667 0.613 13.10% -8.10% 0.563 0.672 0.549 -2.49% -18.30%
Table 6. Performance comparison among ensemble ML-based classifiers with three word embeddings
Classifier Precision Recall F1-score Accuracy
Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B
ML_word2vec 0.776 0.789 0.810 4.38% 2.66% 0.676 0.828 0.697 3.11% -15.82% 0.715 0.808 0.749 4.76% -7.30% 0.735 0.803 0.766 4.22% -4.61%
ML_fastText 0.739 0.777 0.798 7.98% 2.70% 0.669 0.713 0.713 6.58% 0.00% 0.698 0.744 0.753 7.88% 1.21% 0.712 0.754 0.766 7.58% 1.59%
ML_GloVe 0.743 0.759 0.768 3.36% 1.19% 0.656 0.828 0.705 7.47% -14.86% 0.688 0.792 0.735 6.83% -7.20% 0.711 0.783 0.746 4.92% -4.73%
RQ1.1: Performance of Classifiers
For ML-based classifiers, as shown in Table 2, we trained the five types of ML classifiers with three word embedding techniques (i.e., word2vec (w2v), fastText (FT), and GloVe), leading to 15 classifiers (i.e., the five ML models times three word embeddings). We calculated the average values of the classifiers over the three pre-trained word embeddings. For each classifier based on three word embeddings, we underlined the best result of each metric (horizontal comparison); for each metric of the classifiers (precision, recall, F1-score, and accuracy), we marked the best metric result of each classifier on three word embeddings in bold (vertical comparison). From the results, we can see that the kNN classifiers have better scores of Precision than other classifiers with an average Precision of 0.816. As for the remaining metrics, the SVM classifiers achieve relatively better performance on nearly all metrics than other classifiers, with an average Recall of 0.801, F1-score of 0.779, and Accuracy of 0.773. Specifically, the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803.
For DL-based classifiers, as shown in Table 3, we trained one type of classifier, i.e., TextCNN, based on the vocabulary set from our dataset and three word embedding techniques (i.e. in total four classifiers) as mentioned in Section 3.3.1. We can see that TextCNN_FT performs the best compared with the other three DL models.
The ML-based classifiers demonstrated superior performance, achieving, on average, a Precision of 0.753, a Recall of 0.667, an F1-score of 0.700, and an Accuracy of 0.719. In contrast, the DL-based classifiers exhibited lower performance, with an average Precision of 0.574, Recall of 0.664, F1-score of 0.542, and Accuracy of 0.563. These findings suggest that the ML-based classifiers are more effective in accurately classifying the target variable than the DL-based classifiers in our case. Overall, among these classifiers, the SVM classifiers outperform other classifiers, exhibiting the best average results on all metrics, except for the Precision metric.
RQ1.2: Comparison of Embedding Models
From the results in Tables 2 and 3, we can also observe the differences between the classifiers based on three pre-trained word embedding models. As mentioned before, for each classifier based on three word embeddings, we underlined the best result of each classifier on each metric (horizontal comparison). When using the word2vec pre-trained word embedding, most of the ML classifiers have relatively better performance on Precision and Accuracy. Most of the ML classifiers with the fastText model have fairly better F1-scores. Besides, we found that the TextCNN classifier based on the fastText pre-trained word embedding achieves the best performance on all metrics except for Recall, with a Precision of 0.728, a Recall of 0.549, an F1-score of 0.667, and an Accuracy of 0.672.
In addition, we can see in Table 3 that the overall performance of the three TextCNN classifiers with pre-trained word embeddings outperforms the TextCNN classifier based on the vocabulary set from our dataset. This is because pre-trained word embedding models contain more features due to the large textual data used for generating word vectors, and they have been proven to be invaluable for improving the performance in NLP tasks (Qi2018pwe, ; Radford2018ilu, ).
In general, the three pre-trained word embedding models have an impact on the performance of the DL-based classifiers. As shown in Table 3, when we compared their classification performance, in most cases, the classifiers with the fastText model can achieve relatively good performance.
RQ1.3: Comparison of Word Embedding Model Dimensionality
According to the results of RQ1.2, the fastText model can help to generate better results than the other two pre-trained word embedding models. Therefore, to further explore the impact of dimension sizes of the pre-trained word embedding models, we trained the classifiers based on the fastText model and chose three common dimension sizes: 100, 200, and 300. In Table 4, similarly as before, we underlined the best result of each classifier on each metric (horizontal comparison). We can see that five of the six ML/DL classifiers can generate better F1-scores based on the fastText model with 200-dimensional word embedding, compared with 100 and 300-dimensional models.
RQ1 Summary. The SVM classifier trained on the pre-trained word2vec word embedding can achieve the best performance on identifying violation symptoms, with an F1-score of 0.779. In most cases, the classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200 as the dimension size of the pre-trained word embedding models can generate better performance than 100 and 300-dimensional models.
4.2. RQ2: Improving Performance with Voting
To answer RQ2, we employed an ensemble learning technique, i.e., majority voting strategy, to build an ensemble classifier. We compared the performance of individual classifiers with ensemble classifiers. Specifically, we explored to what extent combining one type of classifier based on three word embedding techniques can achieve better performance, as shown in Table 6 and Table 6.
In Table 6, the “Mean” column presents the Average values of each metric of each classifier on three word embeddings from Table 2 and Table 3. The “Best” column denotes the corresponding classifier with the best performance from Table 2 and Table 3. The “Voting” column refers to the performance of the ensemble classifier (based on classifiers with three word embeddings) on each metric after utilizing a majority voting strategy. For example, the Voting values of the SVM classifiers in Table 6 refer to the voting results among the three individual classifiers SVM_w2v, SVM_FT, and SVM_GloVe. “Impro_M” and “Impro_B” represent the improved results of the corresponding ensemble classifiers compared with the results in “Mean” and “Best” columns, respectively.
According to Table 6, when we utilized the voting strategy for each ML classifier based on the three word embeddings, all ensemble classifiers perform better than the mean values of individual classifiers, with F1-score improvement from 0.48% to 13.10%. Table 6 shows that even compared with the best-performing classifiers, most of the ensemble classifiers (4 out of 5) outperform the individual ML/DL classifiers, except for kNN. Moreover, the ensemble TextCNN of the three TextCNN classifiers based on pre-trained word embeddings (see Table 3) does not perform as well as TextCNN_FT (i.e., the best one) on Precision, F1-score, and Accuracy. This means that TextCNN_FT with the fastText pre-trained word embedding performs best compared with the other three TextCNN classifiers, and the voting strategy does not have an obvious improvement for the trained DL classifiers.
Additionally, as shown in Table 6, we utilized the voting strategy for the five ML classifiers based on each of the three word embeddings. For example, ML_word2vec is the ensemble classifier of the five ML classifiers based on the word2vec word embedding (i.e., SVM_w2v, LR_w2v, NB_w2v, DT_w2v, and kNN_w2v). Similarly as before, the “Mean” column refers to the average metric results of ML classifiers (see the last row in Table 2); the “Best” column presents the best-performing classifiers in Table 2; the “Voting” column denotes the performance results of the ensemble classifiers after voting. It is clear that the three ensemble classifiers outperform the average performance results of the individual classifiers, with improvements from 3.11% to 7.98%. Even compared with the best-performing individual classifiers (i.e., the maximum values of the voting samples), the ensemble classifier ML_fastText can still demonstrate performance improvement on all metrics.
In general, regarding the performance difference of ensemble classifiers with three word embeddings, among the selected five ML models, the fastText pre-trained word embedding model achieves relatively better performance improvements on almost all metrics compared with the models using word2vec and GloVe. This finding affirms a similar observation in the result of RQ1.2.
RQ2 Summary. In comparison to individual ML classifiers, the ensemble classifiers achieve better performances after utilizing the majority voting strategy. However, no significant improvement was observed in the performance of the DL classifiers on our dataset when implementing the ensemble strategy.
4.3. RQ3: Validation in Practice
As elaborated in Section 3.4, we designed and conducted a survey and semi-structured interviews to solicit the perceptions from the practitioners on the usefulness of our trained models to automatically identify violation symptoms from code review comments. Note that, our survey and interview did not ask participants to “discover” additional violation symptoms that were not originally discussed in their code reviews.
Overall, we planned to send customized survey invitation emails to 200 code reviewers who were involved in the discussion of the collected review comments concerning architecture violations. However, we could only derive valid email addresses for 169 of them. We finally obtained 24 valid responses to our survey and interviewed 6 participants after sending 169 survey invitation emails and reminder emails; the survey response rate of 14.2%, is a reasonable response rate considering that the general response rate is around 5% in empirical software engineering research (gaese2008CH1, ). Table 7 shows the demographic information of the 24 respondents. Twenty two of the respondents have more than 10 years of experience in software development, while two of them have 6-10 years of experience. Among the respondents, 83.3% (20) are software engineers and 16.7% (4) are architects. Regarding the size of their companies, ten of the respondents reported that their companies have more than 1000 employees, five respondents worked in companies with employees between 501 and 1000, six respondents in companies with employees between 101 and 500, and three respondents in companies with employees between 21 and 100. Besides, the respondents worked in a variety of application domains.
Figure 4. Practitioners’ responses to statements regarding the usefulness of the trained models
Table 7. Demographic information of the survey participants and interviewees
# Country Experience Role Company size Company domain
P1 Ireland More than 10 years Software Engineer employees $\geq$ 1000 HCM services
P2 Germany More than 10 years Software Engineer 21 $\leq$ employees $\leq$ 100 Industrial electronics
P3 Finland More than 10 years Software Engineer 21 $\leq$ employees $\leq$ 100 Medical software
P4 Norway More than 10 years Architect 501 $\leq$ employees $\leq$ 1000 Industrial DataOps
P5 Israel More than 10 years Software Engineer employees $\geq$ 1000 Network
P6 Germany 6 - 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Development tools
P7 USA More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Iaas
P8 Norway More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Cross-platform libraries for application development
P9 Norway More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 SW development frameworks
P10 USA More than 10 years Software Engineer employees $\geq$ 1000 Software development
P11 Germany More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Middleware
P12 USA More than 10 years Software Engineer employees $\geq$ 1000 Research
P13 (I1) Norway More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Development framework
P14 Finland More than 10 years Architect 21 $\leq$ employees $\leq$ 100 Medical devices
P15 Germany More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Tools and frameworks for software development
P16 (I2) Germany More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Middleware Platform
P17 Germany More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Defence development
P18 Denmark 6 - 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Network
P19 (I3) Canada More than 10 years Software Engineer employees $\geq$ 1000 Framework tool
P20 India More than 10 years Architect employees $\geq$ 1000 Development framework
P21 (I4) China More than 10 years Architect employees $\geq$ 1000 Cloud, Data, GenAI
P22 USA More than 10 years Software Engineer employees $\geq$ 1000 Tools for software development
P23 (I5) Germany More than 10 years Software Engineer employees $\geq$ 1000 Game development
P24 (I6) Switzerland More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 AI-related research and development
  • P = Survey Participant, I = Interviewee
Figure 4 presents the feedback from the 24 respondents on the usefulness of the trained models, i.e., the automated identification of violation symptoms from the textual content of code review comments. According to the results, most of the respondents (66.7%) agreed that the violation symptoms identified by our models represent potential architectural violations (Statement 1). 37.5% respondents agreed that the identified violation symptoms can be used to improve the code quality (Statement 2), while 41.7% respondents had a neutral attitude about this statement. 54.2% of the respondents agreed that the identified violation symptoms contain potentially constructive and useful architectural information (Statement 3).
Statements 3 to 6 investigate whether the trained classifiers can play an auxiliary role in development and code reviews. Approximately 41.7% respondents strongly agreed (12.5%) or agreed (29.2%) that the identified violation symptoms can help identify violation-related issues faster than manual identification (Statement 4), while there are still nearly half of the participants convey their uncertainty; the proportion of positive responses indicates that the classifiers may not fully replace manual efforts, but to some extent they can speed up the identification of violation-related issues during code review. Moreover, the majority (62.5%) of the respondents strongly agreed (4.2%) or agreed (58.3%) that automated identification can help locate and prioritize architectural violation issues (Statement 5). Furthermore, 41.7% of the respondents agreed that the identified violation symptoms can inspire them to find out other (or similar) violation symptoms (Statement 6). Notably, one respondent strongly disagreed with all six statements, indicating that he/she might not believe that ML/DL-based classifiers can be useful for identifying architectural violations in practice; we did not treat this response as an outlier. Moreover, only a small number of participants (ranging from 2 for Statement 1 to 5 for Statement 6) chose the “I Don’t Know” option. We further analyzed the responses for the six statements and found no significant statistical differences between the identified violation symptoms. Overall, Figure 4 shows that positive feedback (strongly agree + agree) surpasses negative (disagree + strongly disagree) and neutral feedback in most of the statements, except for Statement 2.
In terms of the semi-structured interviews, the six practitioners that accepted our invitation are denoted as I1 to I6 in Table 7. Each interview lasted approximately 30 to 40 minutes, resulting in a total of 189 minutes of audio recordings. Furthermore, the transcripts of these interviews spanned 35 pages with over 22K words. In general, all of them expressed a positive attitude towards such an automated way to identify violation symptoms from code review comments. We present the key common points of the interviewees’ opinions below:
  • (IQ1) Potential benefits: The integration of such an automated identification of violation symptoms of architecture erosion could help to reduce the time and effort of developers, which is recognized by four interviewees. As I3 stated, “extra checking could make our work faster and facilitate the code review workflow”. Additionally, I4 mentioned that “Code reviewers prioritize two main categories of issues: those related to functionality and those concerning the architecture, and such an approach is particularly valuable and beneficial for identifying and addressing architectural problems”.
  • (IQ1) Challenges: Three out of six interviewees raised concerns regarding data leakage problems when employing or integrating such an automated approach supported by third-party libraries or specific LLMs, especially for inner-source projects. Besides, two out of six interviewees expressed their concerns about “Crying Wolf” effect of false positives; as I1 mentioned: “Any tool that gives you warnings that are false, after a while you start ignoring its warnings”. If tools give false warnings (or warnings that users consider trivial), it eventually leads users to ignore all warnings, including legitimate ones. This calls for the interpretability of such automated approaches that participants might not only want to know about the identified violation symptoms, but also understand why they are flagged as “violations”; as I2 mentioned: “As a user, I am not satisfied to be told where the issues are located, and I also want to know why they are flagged as violations with explanations”. Therefore, putting more effort into improving the interpretability could pose another challenge for such automated approaches.
  • (IQ2) Strategies: I4 thought that the best practices and specific rules (e.g., guidelines on architectural layering, module dependencies, and coupling thresholds) for identifying typical violation symptoms (e.g., layering violations) should be considered and integrated when employing automated techniques during code review. This could help customize such an approach of identifying violation symptoms and then maximize the identification accuracy for inner-source projects inside companies.
  • (IQ3) Visions for improving code quality: With the advent of LLMs, five out of six interviewees are willing to utilize our approach with customized LLMs for specific tasks in the future, e.g., identifying architecture issues by feeding architecture constraints to the LLMs and providing architecture refactoring suggestions.
RQ3 Summary. Practitioners’ responses indicate that the automated identification of violation symptoms is perceived as useful in practice. In particular, it helps practitioners identify potential violation symptoms earlier, better recognize architecturally related violations and prioritize violation-related issues during review and maintenance activities.
4.4. RQ4: Comparison with LLM-based Classifiers
To answer RQ4, we conducted additional comparative experiments by constructing LLM-based classifiers. As shown in Table 8, we evaluated three LLMs across five shot settings (one zero-shot and four few-shot configurations). Qwen-2.5 demonstrates the highest Precision, achieving a peak of 0.873 in the 4-shot setting and the highest average Precision of 0.839. In contrast, GPT-4o excels in Recall, with a maximum value of 0.951 at 10-shot and an average Recall of 0.879. DeepSeek-R1 also performs well in Recall, particularly at 4-shot (0.943), resulting in the highest average Recall of 0.898. For F1-scores, GPT-4o achieves the highest peak 0.851 at 2-shot and the highest average 0.811, followed by Qwen-2.5 (0.782 at 6-shot) and DeepSeek-R1 (0.767 at 4-shot). GPT-4o also leads in Accuracy, reaching a peak of 0.848 at 2-shot and averaging 0.793. Qwen-2.5 achieves its highest Accuracy (0.799) at 6-shot, while DeepSeek-R1 has the lowest average Accuracy (0.696) despite a 2-shot peak of 0.713.
Overall, GPT-4o provides the most balanced and robust performance across all metrics and shot settings. Qwen-2.5 leads in Precision and maintained competitive Accuracy, while DeepSeek-R1 stands out in recall but shows lower Precision and overall Accuracy.
Table 8. Performance comparison of different LLMs with varying shot settings
Model Precision Recall
0-shot 2-shot 4-shot 6-shot 10-shot Average 0-shot 2-shot 4-shot 6-shot 10-shot Average
Qwen-2.5 0.825 0.831 0.873 0.854 0.811 0.839 0.656 0.566 0.566 0.721 0.705 0.643
GPT-4o 0.701 0.835 0.851 0.699 0.712 0.759 0.844 0.869 0.795 0.934 0.951 0.879
Deepseek-R1 0.653 0.657 0.646 0.611 0.635 0.640 0.803 0.893 0.943 0.926 0.926 0.898
Average 0.726 0.774 0.790 0.721 0.719 0.746 0.768 0.776 0.768 0.860 0.861 0.807
Model F1-score Accuracy
0-shot 2-shot 4-shot 6-shot 10-shot Average 0-shot 2-shot 4-shot 6-shot 10-shot Average
Qwen-2.5 0.731 0.673 0.687 0.782 0.754 0.725 0.758 0.725 0.742 0.799 0.771 0.759
GPT-4o 0.766 0.851 0.822 0.800 0.814 0.811 0.742 0.848 0.828 0.766 0.783 0.793
Deepseek-R1 0.721 0.757 0.767 0.736 0.753 0.747 0.689 0.713 0.713 0.668 0.697 0.696
Average 0.739 0.760 0.759 0.773 0.774 0.761 0.730 0.762 0.761 0.744 0.750 0.749
Compared to ML/DL-based classifiers (see Tables 2 and  3), the LLM-based classifiers achieve better performance on the same classification task, with average Precision of 0.746, Recall of 0.807, F1-score of 0.761, and Accuracy of 0.749. This finding indicates that the LLM-based classifiers are more effective in identifying the violation symptoms than traditional ML/DL-based classifiers. Among the LLM-based classifiers, the GPT-4o model with a 2-shot setting outperforms others in average performance across all metrics except Recall.
We also examined whether ensemble strategy across different shot settings could improve performance (see Tables 10 and 10). Table 10 shows the results of ensemble classifiers based on a voting strategy. The “Mean” column refers to the average performance of each metric across all LLMs and shot settings from Table 8, and the “Best” column captures the best performance per shot setting. The “Voting” column reports the ensemble performance obtained via majority voting among models with the same shot setting. Likewise, the improvements over the “Mean” and “Best” are shown in the “Impro_M” and “Impro_B” columns, respectively. We observed that the ensemble classifiers outperform the average results of different shot settings across all metrics in Table 10. However, they generally underperform when compared to the best individual classifiers in Precision and Recall. In contrast, the ensemble classifiers slightly improve over the best individual classifiers in F1-score and Accuracy.
Similarly, Table 10 examines ensemble classifiers within a single LLM across different shot settings. The “Mean” column presents the average metric values across all five shot settings from Table 8; the “Best” column shows the best performance of single LLMs; the “Voting” column displays the results from the ensemble classifiers across shot settings using a voting strategy. Table 10 shows that, for GPT-4o and Qwen-2.5, the ensemble classifiers outperform their respective average performances but do not exceed their best individual results. Notably, for Deepseek-R1, the ensemble classifier surpasses both the average and best individual classifiers across all metrics except Precision.
Table 9. Performance comparison of ensemble LLM-based classifiers with different shot settings
Classifier Precision Recall F1-score Accuracy
Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B
0-shot 0.726 0.825 0.760 4.65% -7.85% 0.768 0.844 0.779 1.42% -7.77% 0.739 0.766 0.769 4.09% 0.45% 0.730 0.758 0.766 5.06% 1.08%
2-shot 0.774 0.835 0.813 5.01% -2.59% 0.776 0.893 0.820 5.63% -8.26% 0.761 0.851 0.816 7.34% -4.12% 0.762 0.848 0.816 6.99% -3.86%
4-shot 0.790 0.873 0.845 6.92% -3.27% 0.768 0.943 0.803 4.63% -14.78% 0.758 0.822 0.824 8.58% 0.18% 0.761 0.828 0.828 8.80% 0.00%
6-shot 0.722 0.854 0.717 -0.61% -16.07% 0.861 0.934 0.893 3.81% -4.39% 0.773 0.800 0.796 2.95% -0.55% 0.745 0.799 0.770 3.49% -3.59%
10-shot 0.719 0.811 0.723 0.46% -10.94% 0.861 0.951 0.918 6.67% -3.45% 0.774 0.814 0.809 4.49% -0.66% 0.750 0.783 0.783 4.37% 0.00%
Table 10. Performance comparison among ensemble LLM-based classifiers
Classifier Precision Recall F1-score Accuracy
Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B Mean Best Voting Imp_M Imp_B
Qwen-2.5 0.839 0.873 0.849 1.24% -2.74% 0.643 0.721 0.648 0.77% -10.23% 0.725 0.782 0.735 1.31% -6.05% 0.759 0.799 0.766 0.97% -4.10%
GPT-4o 0.759 0.851 0.803 5.72% -5.64% 0.879 0.951 0.902 2.61% -5.17% 0.811 0.851 0.849 4.78% -0.23% 0.793 0.848 0.840 5.89% -0.97%
Deepseek-R1 0.640 0.657 0.650 1.47% -1.05% 0.898 0.943 0.943 4.93% 0.00% 0.747 0.767 0.769 3.01% 0.33% 0.696 0.713 0.717 3.06% 0.57%
RQ4 Summary. GPT-4o shows the most balanced performance, leading in F1-score and Accuracy, while Qwen-2.5 excels in Precision and DeepSeek-R1 in Recall. LLM-based classifiers outperform traditional ML/DL-based classifiers across all metrics. Ensemble classifiers offer slight improvements in F1-score and Accuracy but do not surpass the best individual LLM-based classifiers in Precision and Recall.
5. Discussion
5.1. Interpretation of Results
5.1.1. RQ1: Identifying Violation Symptoms
Our trained ML/DL classifiers can identify violation symptoms in code review comments. The results of RQ1.1 indicate that certain factors such as the type of classifier, the pre-trained word embeddings, and hyperparameters influence the performance of classifiers. Specifically, we found that the SVM classifier with word2vec outperforms other ML/DL-based classifiers. We also found that DL classifiers do not always have better performance than ML classifiers, and the outcome may stem from factors such as model complexity, dataset size, and hyperparameter settings; we note that training DL classifiers takes much longer time than training ML classifiers, as DL classifiers have much more complex network architecture and more parameters. In recent years, some studies that leveraged DL-based techniques for SE tasks have achieved better performance that all other state-of-the-art techniques (e.g., ML-based techniques) (Yang2022pmse, ). However, the achieved results imply that ML classifiers might be enough for certain text classification tasks with small-size datasets, especially as the trained DL classifiers do not show performance improvement in such a scenario. This finding is also observed in previous studies (Liu2018nmt, ; Fu2017eoh, ): for certain SE tasks, simple ML-based approaches are capable of achieving equal or even better performance than DL-based approaches within less time. Thus, we advise researchers to not blindly apply DL-based techniques and consider simple approaches first.
Finding 1: Traditional ML-based approaches have sufficient performance for many SE tasks, such as identification of architecture violations in code reviews.
From the results of RQ1.2, we can see that the TextCNN classifier based on the fastText pre-trained word embedding can perform better than other DL models. We thus emphasize that the fastText pre-trained technique has a better capacity for training well-performed DL models; this also aligns with the finding in the work of Sesari et al. (Sesari2022pte, ), that is, the fastText pre-trained word embedding model has less bias compared to word2vec and GloVe.
Besides, the results of RQ1.3 show that 200-dimensional pre-trained word embedding models outperform 100 and 300-dimensional models. The possible reason is that 200-dimensional embeddings strike a balance between expressiveness and robustness for our dataset size: lower-dimensional (100d) embeddings under-represent nuanced lexical patterns, whereas higher-dimensional (300d) embeddings introduce excess parameters, increasing the risk of overfitting. Our results are slightly different from previous studies (e.g., (Ren2019nndSATD, ; Li2022SATD, )) where the 300-dimensional word embedding model has better performance. One possible reason is that the previous studies only measured the average results of classifiers with different dimensional word embedding models, and the average results might be influenced by the individual results. Our results present and compare the classification performance of each ML/DL classifier, which could be more accurate.
Finding 2: The 200-dimensional fastText pre-trained word embedding model can be used to train classifiers with better performance when conducting binary classification of violation symptoms from code review comments.
In general, due to a large number of hyperparameters of DL classifiers, we did not specifically focus on the best combination of hyperparameters of the TextCNN classifier, but certain key hyperparameters such as the dimension size of word embedding model. Thus, future work can consider performing hyperparameter tuning (e.g., utilizing the grid search technique) to explore other possible combinations of the hyperparameters to further improve performance.
5.1.2. RQ2: Improving Performance with Voting
Evaluating the performance of an ensemble classifier offers insight into the performance gains compared with individual classifiers. For voting-based approaches, an accurate output is attained only when a majority of the constituent models generate the correct output (hansen1990nne, ). Our results provide further empirical evidence for the efficacy of the majority voting strategy on improving the text classification performance. This technique effectively combines the predictions of multiple classifiers to reach a final decision, leading to improved overall classification performance. Nevertheless, we observed that such performance improvement is not obvious in TextCNN in our case. One possible reason is that our DL classifiers (all TextCNN variants) and the modest dataset size prevented the majority‑vote ensemble from yielding significant gains. We argue that the voting strategy remains valuable to minimize potential prediction errors if there are more outputs for voting.
Finding 3: Compared with individual ML/DL-based classifiers for identifying architecture violations from code review comments, ensemble learning is an effective way for improving classification performance. In most cases, ensemble classifiers that utilize the voting strategy can enhance all classification performance metrics.
5.1.3. RQ3: Validation in Practice
The existence of architectural violations can be challenging, as they can impede practitioners’ understanding of architecture. This particularly influences novice developers who lack practical familiarity with architecture. The practitioners we surveyed and interviewed recognized the identified issues in code snippets, as they had previously noted them during code review. However, they may not have realized these issues were violation symptoms of architecture erosion. Our approach helps practitioners realize the existence of violation symptoms of architecture erosion. The results of RQ3 indicate a predominantly positive feedback, confirming the usefulness of our trained ML/DL classifiers in potentially helping developers during maintenance and evolution. Through analyzing the received responses to the statements (especially Statements 3, 4, and 5), we found that developers convey interest in utilizing automated approaches, tools, or plug-ins to identify violation symptoms during code review, recognizing their potential to enhance the productivity of dealing with architectural violations. Overall, the survey and interview results indicate that developers hold positive attitudes towards such kinds of tools. To some extent, their perception might be influenced by the use of balanced datasets in classifier evaluation. Future work could focus on deploying the classifiers in more real-world scenarios, where data distributions may differ and classifiers’ performance could vary, potentially improving or declining compared to our current results. Moreover, automated techniques regarding identifying potential architecture violations could be supplementary means to minimize architecture violations and improve code quality during code review. For example, such techniques could provide code reviewers with warnings that architecture violations need to be refactored, based on the review discussions. Furthermore, according to the participants’ feedback, automated techniques could help code reviewers prioritize such refactorings based on various criteria (e.g., the importance of involved components, the number of violation symptoms in those components, and the seniority of code reviewers). Such a prioritization could help to fix violation symptoms that might otherwise have been ignored.
Moreover, according to the respondents’ feedback, certain practitioners might hold a conservative attitude toward ML/DL-based classifiers for identifying violation symptoms, which could be observed from the “Neutral” results in Figure 4, especially for Statement 2. One reason is that they might be skeptical about the performance of the classifiers (false positives or negatives), casting doubt about the reliability of identified violation symptoms. Nevertheless, no classifier can ever achieve 100% accuracy in classification tasks, and there is always a trade-off between precision and recall. This disclaimer should be made clear when showing classification tasks to practitioners. Another potential reason could be that certain identified review comments may be part of the architectural design discussions during the development process, rather than the final architectural decisions. This is particularly common in agile development, where architectures evolve over time and certain decisions are continuously discussed and refined. Such concerns are beyond the target of our research, as the trained classifiers can only provide judgments based on the collected violations, instead of exploring the evolution of architecture violations.
Finding 4: Despite a few concerns about the accuracy of the identified violation symptoms, the vast majority reacted favorably, indicating strong practitioner support for automated identification of violation symptoms during code review.
5.1.4. RQ4: Comparison: ML/DL-based vs. LLM-based Classifiers
The results of RQ4 clearly demonstrate that LLMs when configured as classifiers via prompt engineering, substantially outperform traditional ML/DL-based approaches on the violation symptom identification task. Across all metrics, the average performance of LLM-based classifiers exceeds that of both ML/DL-based classifiers. This suggests that LLMs’ deep contextual understanding and few-shot adaptability confer distinct advantages for text classification tasks.
Among the three LLM-based classifiers, the results show that different LLMs vary in performance on the automated identification of violation symptoms. The possible reason could be that different LLMs draw on distinct training corpora and pre-training goals, which in turn shape their strengths and weaknesses across SE tasks, such as program repair, code generation. Consequently, practitioners should align task requirements (e.g., code understanding, text classification) with each LLM’s underlying training focus to select suitable LLMs.
Furthermore, ensemble strategy either across models (see Table 10) or across shot settings for an individual model (see Table 10) yields modest but consistent improvements over average results of various shot settings from Table 8. However, ensemble classifiers generally do not outperform the best-performing individual LLMs in Precision or Recall, underscoring that an ensemble’s strength lies in balancing errors rather than maximizing peak performance capabilities. Notably, employing a voting strategy cannot significantly improve the performance of ensemble LLM-based classifiers. A plausible explanation could be that the individual LLM-based classifiers already achieve very good performance, leaving limited room for further enhancement through ensembling. In fact, the ensemble technique may even degrade performance, potentially due to the amplification of inherent issues such as LLM-induced hallucination (GPT4, ). This phenomenon closely resembles the Law of Diminishing Marginal Utility in economics, whereby additional inputs yield progressively smaller benefits beyond a certain threshold (jevons2013PE, ).
Finding 5: LLM-based classifiers typically outperform traditional ML/DL-based classifiers for identifying violation symptoms. However, ensembling multiple LLM-based classifiers does not yield further performance gains compared to individual LLM-based classifiers, potentially due to the compounding effect of hallucinations inherent in LLMs.
5.2. Implications
Establish and enrich datasets. Many studies utilized or proposed automated approaches to identify or predict knowledge-related information from textual artifacts. However, no studies focused on the automated identification of violation symptoms from code review comments. This suggests that there is a need for more empirical reports and follow-up tools regarding such issues. Our study provides a starting point for researchers to automatically identify architecture violations in textual artifacts such as review comments. On top of this, we encourage researchers and practitioners to establish and share datasets related to architecture violations to lay a solid foundation for future research.
Moreover, our study is scalable in practice, as practitioners can retrain superior classifiers by adding more code review comments of violation symptoms from different projects, either manually or by using our classifiers. Besides, mining architecture violations from multiple textual artifacts (e.g., issues, source code comments, and pull requests) can provide a more comprehensive examination of architectural violations and also enrich the dataset used for identifying such violations. Furthermore, our findings regarding the experiments in this study might benefit or be adapted by researchers and practitioners to identify other issues, such as security and vulnerability discussions.
Suggestion 1: Researchers and practitioners are encouraged to establish and share datasets related to architecture violations. Besides, our study is scalable to retrain superior classifiers by adding more data from different sources, and can also be adapted to identify other issues.
Choose appropriate classifiers. For traditional ML/DL-based classifiers, an ensemble classifier is a collection of classifiers that work together to classify instances by combining their individual outputs through voting. Researchers can consider exploring different selections of classifiers to further enhance the performance of ensemble classifiers. For example, to some extent, assigning varying weights to individual classifiers might partially mitigate the impact of biased classifiers on voting results. While ensemble classifiers based on LLMs require more time and resources, their performance often does not exceed that of certain top-performing individual LLM-based classifiers, particularly when configured with optimal few-shot settings.
Suggestion 2: When constructing traditional ML/DL-based classifiers, combining classifiers and assigning different weights to individual classifiers are recommended to further enhance the classification performance of ensemble classifiers for identifying architecture violations. In contrast, for LLM-based classifiers, choosing a suitable few-shot setting is recommended over ensemble techniques.
Optimize code review practice. In the future, researchers and practitioners can explore the development of plug-ins for code review platforms (e.g., Gerrit) that provide warnings of violations to developers or even suggestions to fix those violations. Such plug-ins can be developed utilizing historical code review data or harnessing the existing LLMs through fine-tuning. These plug-ins have the potential to assist developers and maintainers in conducting comprehensive examinations of preexisting architectural violations, and may even serve as reminders of their presence. Additionally, implementing the trained classifiers (e.g., as an extension) into the existing tools (or toolsets) is also worthy of consideration in practice. For example, code review platforms can integrate such classifiers to empower the automated code review process. Such tools may provide more insights and help novices or maintainers troubleshoot architecturally relevant violation issues.
Suggestion 3: Researchers and practitioners can consider integrating classifiers as plug-ins for code review platforms (e.g., Gerrit) to assist developers in identifying potential architecture violations, or integrating customized classifiers into existing tools to optimize the automated code review process.
6. Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats by following the guidelines proposed by Wohlin et al(Wohlin2012ESE, ). Internal validity is not considered, since this study does not address any causal relationships between variables.
6.1. Construct Validity
Construct validity concerns the connection between operational measures and the studied subjects. The first potential threat to construct validity is the evaluation of our selected ML/DL-based and LLM-based classifiers. Regarding the classifiers’ performance evaluation, we used the four metrics (recall, precision, F1-score, and accuracy) that are widely used to evaluate the performance of various automated software engineering techniques (see Section 7.3). Therefore, the metrics pose little threat to the construct validity. Another potential threat comes from the convenience sampling method of the survey and interview, as individuals with negative views towards the research might be less inclined to participate in our survey and interview. This is a compromise inherent to the convenience sampling method, and it implies that this threat cannot be completely mitigated. Nevertheless, to improve the response rate and collect as many opinions as possible from practitioners, we sent two rounds of reminder emails to the involved practitioners. It is also possible that some survey respondents did not understand the statements well. To reduce this threat to construct validity, we conducted a pilot survey for refining the survey design. Besides, we provided an “I Don’t Know” option in our survey. Moreover, we also interviewed six participants to collect in-depth opinions besides the six defined statements in the survey. To mitigate this threat, we offered an optional open-ended question that allows respondents to freely share their thoughts with us.
6.2. External Validity
External validity pertains to the generalizability of our experiment results and findings. In certain situations, textual artifacts (e.g., code review, architecture discussions) may be unavailable. In such scenarios, we recognize that our approach may be complemented with traditional code-based architecture conformance checking techniques. Exploring the integration of such an automated approach into existing IDEs or tools remains part of our future work. As noted by Peter et al. (peters2017text, ), it is not possible to use an exhaustive approach to test all learning models. Therefore, we selected five common ML algorithms, one DL algorithm, three popular pre-trained word embedding models, and three well-performed LLMs, as they are all widely used to conduct text classification. We admit that using different algorithms, pre-trained word embedding models, and LLMs may generate varying results, despite their popularity. We cannot completely mitigate this threat. Selecting different datasets and experimental settings might influence the generalizability of (part of) our findings. We acknowledge that the chosen open-source projects are not representative of all projects from the OpenStack and Qt platforms, and our findings may not be generalized to diverse projects, particularly those from different domains such as closed-source projects or projects on other platforms like Apache. Furthermore, the participants of the survey and interviews are contributors to the selected projects, therefore, the findings of RQ3 might not be generalized to other projects. To support further generalizability, we have made our dataset and scripts (onlinepackage, ) available to enable other researchers and practitioners to replicate, reuse, and extend our study.
6.3. Reliability
Reliability reflects the replicability of a study regarding yielding the same or similar results when other researchers reproduce this study. One potential threat to reliability comes from the experimental settings (e.g., input data and parameters for training classifiers) used in this work. To address this, we used the dataset from our previous study (Li2023wvs, ) that had publicly shared the labeled data and corresponding scripts for collecting data. Besides, we also shared the source code in our replication package (onlinepackage, ) to facilitate reproduction and validation. Another threat might concern the replicability of the survey and interview. In the current study, biases related to survey design were mitigated through careful question formulation and pilot testing. While such biases cannot be entirely eliminated, we provided the survey template, semi-structured interviews, and customized emails in our replication package (onlinepackage, ) to maximize transparency and reliability, in order to facilitate replication in future studies.
7. Related Work
In this section, we briefly discuss the current state of research in three associated areas of our study. We first describe related work on code review comments in Section 7.1; we then introduce previous studies on architecture violations in Section 7.2; we finally present the benefits of utilizing ML/DL and LLM-related techniques to analyze artifacts in software development in Section 7.3.
7.1. Code Review Comments
Code review comments contain massive knowledge related to software development, and a variety of studies analyze software defects and evolution through mining review comments and commit records. Han et al. (Han2022cmd, ) conducted an empirical study by manually checking 25,415 code review comments to investigate code smells identified in modern code review from four OSS projects. Their results show that most reviewers provide constructive suggestions to help developers fix code, and developers are willing to repair the smells through suggested refactoring operations. El Asri et al. (ElAsri2019ess, ) empirically investigated the impact of expressed sentiments on the code review duration and its outcome. They found that reviews with negative comments on average took longer time to complete than the reviews with positive/neutral comments. Kashiwa et al. (Kashiwa2022SATDmcr, ) reported Self-Admitted Technical Debt (SATD) during code review by checking 156,372 review records from OpenStack and Qt. They found that 28-48% SATD are introduced during code review, and 20% SATD comments are created because of reviewers’ requests.
Besides, Uchôa et al. (Uchoa2020hdm, ) investigated the impact of code review on the evolution of design degradation through mining and analyzing a plethora of code reviews from seven OSS projects. They found that there is a wide fluctuation of design degradation during the revisions of certain code reviews. Paixão et al. (Paixao2020bti, ) explored how developers perform refactorings in code review, and they found that refactoring operations are most often used in code reviews that implement new features. Besides, they observed that the refactoring operations were rarely refined or undone along the code review, and such refactorings often contribute to new code smells and bugs.
7.2. Architecture Violations
Over the past decades, there have been various investigations on architecture violations. Brunet et al. (Brunet2012av, ) performed a longitudinal study to explore the evolution of architectural violations in 19 bi-weekly versions of four open source systems. They investigated the life cycle and location of architecture violations over time by comparing the intended and recovered architectures of a system. They found that architectural violations tend to intensify as software evolves and a few design entities are responsible for the majority of violations. More interestingly, some violations seem to be recurring after being eliminated. Mendoza et al. (Mendoza2021avd, ) proposed a tool called ArchVID based on model-driven engineering techniques for identifying architecture violations from source code; this tool supports recovering and visualizing the implemented architecture.
Besides, Terra et al(Terra2015rsr, ) reported their experience in fixing architectural violations. They proposed a recommendation system ArchFix that provides refactoring guidelines for developers and maintainers to repair architectural violations in the module architecture view of object-oriented systems. Their results show that their approach can provide correct recommendations for 75% and 79% of the detected architecture violations on two selected systems, respectively. Maffort et al(Maffort2016mav, ) proposed an approach to check architecture conformance for detecting architecture violations based on defined heuristics. They claimed that their approach relies on the defined heuristic rules and can rapidly raise architecture violation warnings.
7.3. Analyzing Software Repositories with ML/DL and LLM Techniques
The application of ML/DL and LLM techniques to extract informative knowledge from various artifacts during software development is gaining popularity in the SE community. In recent years, a plethora of empirical studies have been conducted to investigate automated techniques for supporting SE tasks. Khalajzadeh et al. (Khalajzadeh2022sda, ) conducted an empirical study by extracting and manually analyzing app reviews from 12 OSS projects and developed ML/DL models to automatically detect and classify human-centric issues from app reviews and discussions. Their results show that automated techniques can help developers to recognize and appreciate human-centric issues of end-users more easily. Nasab et al. (Nasab2021ais, ) developed 15 ML/DL models to automatically identify security discussions, and then they collected practitioners’ feedback through a validation survey and confirmed the promising applications of the models in practice.
Hou et al. (hou2024large, ) conducted a systematic literature review to explore the applications of LLMs across various phases of software development, from requirements gathering to program repair. Their study provides a comprehensive overview of recent advancements in the applications of LLMs in software engineering tasks, highlighting significant opportunities and critical areas that require further research and development to fully leverage LLMs in software engineering. Ahmed et al. (Ahmed2025can, ) investigated the potential of LLMs to substitute costly human annotation efforts in evaluating SE artifacts. They found that state-of-the-art LLMs can achieve inter-rater agreements comparable to human annotators across various code and code-related tasks. De Martino (Martino2025LLM4MSR, ) et al. presented a mixed-method empirical study combining a rapid literature review and survey to examine the use of LLMs in Mining Software Repositories (MSR). They uncovered 14 approaches and actions to mitigate 9 identified threats in applying LLMs to MSR, and proposed PRIMES 2.0, an enriched six-stage empirical framework with 23 substeps to enhance transparency, reproducibility, and methodological robustness in LLM-based MSR studies.
7.4. Conclusive Summary
Table 11 summarizes the key differences between our work and prior studies. While previous studies addressed relevant topics pertaining to code review, architecture violations, and the use of ML/DL/LLM techniques in mining software repositories, none of them have specifically utilized these automated approaches to classify architecture violations during code review. Building on developers’ perspectives, we further collected feedback from practitioners involved in the code review discussions to evaluate the usefulness of the classifiers. In contrast to previous studies listed in Table 11, our work not only complements prior studies by bridging the gap between identifying violation symptoms and applying automated techniques, but also incorporates practitioners’ feedback to evaluate these automated techniques.
Table 11. Comparison between our work and previous studies
Our Work Previous Studies Comparison
Automatically identifying violation symptoms in code review comments. We explored the application of automated techniques to specifically identify violation-related issues (e.g., violation symptoms) from code review comments. Previous studies (Han2022cmd, ), (ElAsri2019ess, ), (Kashiwa2022SATDmcr, ), (Uchoa2020hdm, ), (Paixao2020bti, ) focused on the code-level issues and typically involved manual review during the code review process. Unlike previous empirical studies that manually analyzed code review comments, we attempted to employ automated techniques to identify architecture-level issues (specifically violation symptoms) and compared the performance differences among different automated techniques.
Investigating architecture violations during code review. Discussions regarding architecture violations can occur in the code review process, and we specifically investigated violation symptoms from textual artifacts like code review comments. Previous studies (Mendoza2021avd, ), (Brunet2012av, ), (Terra2015rsr, ), (Maffort2016mav, ) typically analyzed architecture violations through longitudinal study, tools, and defined rules, and often inspected architecture violations in source code during development or maintenance phases. Compared to previous studies focusing on architecture violations from source code, we proposed an automated approach to identify architecture violations during code review, and our approach can act as a complementary way to modern code review.
Exploring the application of automated techniques to identify violation symptoms. We investigated the feasibility of automated techniques for analyzing violation symptoms by employing various configurations of the trained classifiers. Moreover, we evaluated the usefulness of our approach by collecting practitioners’ feedback. Previous studies (hou2024large, ), (Nasab2021ais, ), (Khalajzadeh2022sda, ), (Ahmed2025can, ), (Martino2025LLM4MSR, ) utilized different automated approaches (such as ML, DL, and LLM techniques) in various SE tasks, including identifying security issues, classifying human-centric app reviews. Differentiated from previous studies utilizing automated techniques to diverse SE tasks, our work specifically explored the feasibility of training traditional ML-based classifiers, DL-based classifiers, as well as state-of-the-art LLM-based classifiers to identify violation symptoms in code review comments. We compared the performance of these models across different configurations to provide more empirical insights. We also solicited practitioners’ viewpoints on the usefulness of our approach.
8. Conclusions and Future Work
During code review, reviewers typically spend a substantial amount of effort in comprehending code changes, as significant information (e.g., architecturally relevant information) for inspecting code changes may be dispersed across several files that the reviewers are not acquainted with. Automated identification of architecture violations from the discussions between reviewers and developers can save valuable time and effort to locate and check potential issues.
In this work, we attempted to address the issue of automated identification of violation symptoms in code review comments. To this end, we performed a series of experiments on a dataset of architecture violations from code review comments. Subsequently, to validate the usefulness of our trained ML/DL-based classifiers, we conducted a survey and semi-structured interviews that acquired feedback from the involved developers who discussed architecture violations in code reviews. Furthermore, to compare LLM-based classifiers with traditional ML/DL-based classifiers, we performed the same classification task using three state-of-the-art LLMs (i.e., GPT-4o, Qwen-2.5, and DeepSeek-R1).
Specifically, we developed 15 ML-based and 4 DL-based classifiers to identify violation symptoms from developer discussions in code reviews (i.e., code review comments from four OSS projects in Gerrit). The results show that the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100 and 300-dimensional models. In addition, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperforms the individual classifiers. Moreover, practitioners’ perception of the usefulness of the classifiers’ results validates the promising applications of automated identification of violation symptoms in practice. Especially, such classifiers can enable and inspire practitioners to find architecturally-related violation issues, prioritize, and eventually handle them. Additionally, LLM-based classifiers indeed outperform traditional ML/DL-based classifiers on identifying violation symptoms, and GPT-4o achieves comparatively better performance than Qwen-2.5 and DeepSeek-R1. Interestingly, we found that the ensemble technique cannot further improve the performance of LLM-based classifiers.
In the future, we plan to conduct a more in-depth investigation of how practitioners deal with architecture violations in code review, what kind of practices and tools they use in handling architecture violations, and evaluate whether our approach can support the identification of architecture violations in their daily job (e.g., through controlled experiments to compare the effectiveness of our approach with human-centric practices). Moreover, we want to evolve the classifiers for identifying fine-grained types of architecture violations based on embeddings for code review comments and a larger dataset of more projects from both open-source and industry projects. In particular, a bigger picture is to delve into a more in-depth investigation of the performance of trained classifiers as plug-ins or tools using more real-world scenarios. The goal of such plug-ins or tools for code review platforms is to provide warnings of potential architecture violations to practitioners (e.g., developers and reviewers) for them to re-check the related code changes during code review.
Data Availability
The replication package of this work has been made available at (onlinepackage, ).